### v2

# SNOOPI: Supercharged One-step Diffusion Distillation with Proper Guidance

Abstract
Recent approaches have yielded promising results in distilling multi-step text-to-image diffusion models into one-step ones. The state-of-the-art efficient distillation technique, i.e., SwiftBrushv2 (SBv2), even surpasses the teacher model’s performance with limited resources. However, our study reveals its instability when handling different diffusion model backbones due to using a fixed guidance scale within the Variational Score Distillation (VSD) loss. Another weakness of the existing one-step diffusion models is the missing support for negative prompt guidance, which is crucial in practical image generation. This paper presents SNOOPI, a novel framework designed to address these limitations by enhancing the guidance in one-step diffusion models during both training and inference. First, we effectively enhance training stability through Proper Guidance - SwiftBrush (PG-SB), which employs a random-scale classifier-free guidance approach. By varying the guidance scale of both teacher models, we broaden their output distributions, resulting in a more robust VSD loss that enables SB to perform effectively across diverse backbones while maintaining competitive performance. Second, we propose a training-free method called Negative-Away Steer Attention (NASA), which integrates negative prompts into one-step diffusion models via cross-attention to suppress undesired elements in generated images. Our experimental results show that our proposed methods significantly improve baseline models across various metrics. Remarkably, we achieve an HPSv2 score of 31.08, setting a new state-of-the-art benchmark for one-step diffusion models.
1 Introduction
Diffusion models have recently become popular due to their capacity to produce high-quality, diverse outputs in image [34, 39, 2, 52, 4], 3D [35, 48, 23], audio [17, 10] and video [53, 50, 3, 11] synthesis. Unlike other generative models, such as Generative Adversarial Networks (GANs), diffusion models gradually refine their output through a series of steps, allowing them to achieve diverse yet impressively detailed and realistic outputs. However, this iterative refinement is both time-consuming and computationally demanding, which limits their practicality in real-world applications. As a result, there’s increasing interest in techniques to speed up diffusion models while maintaining output quality. Several strategies are being explored to enhance their efficiency, including optimizing model architectures, improving sampling methods, and using latent space representations instead of image space. Recently, a particularly promising approach has gained traction: timestep distillation. Researchers have found that by adding specific constraints while training another model to mimic a pre-trained multi-step teacher, they can reduce the required steps, creating a faster and more efficient student that retains much of the original quality with minimal adjustments.
Figure 1: Comparison of various methods, including SNOOPI and alternatives (SwiftBrush, DMD, YOSO, PG-SB), using the PixArt-$\alpha$ backbone. Each method generates an image with the prompt ”Vincent van Gogh” (green). PG-SB helps stabilize the training process of image-free distillation, addressing the typical instability in such training. SNOOPI, a combination of PG-SB and the NASA module, enables effective negative prompting (red) even for one and few-step models, allowing for the removal of specific unwanted features and providing enhanced control over image attributes.
Current approaches for accelerating text-to-image diffusion models fall into two main categories: image-dependent training and image-free training methods. Image-dependent methods, which rely on real images or synthetic ones generated by the teacher models, include adversarial-based approaches [41, 19, 42, 51] and reconstruction-based methods [28, 56, 12, 19]. While effective, these methods demand huge storage, high memory, extensive computational power and might suffer from mode collapse due to too strong image supervision. In contrast, image-free methods [33], which mainly use score distillation losses like Score Distillation Sampling (SDS) [36] or Variational Score Distillation (VSD) [48], avoid the need for image datasets, instead requiring minimal data, often just text prompts. Models such as SwiftBrushv2 (SBv2) [7], even when omitting image-regularization term, making it a fully image-free distillation approach, have achieved impressive quality, even outperforming their multi-step counterparts on certain benchmarks. Hence, in this paper, we focus on the image-free approach and start with SBv2 as the baseline.
However, our research identifies two key challenges in current one-step distillation approaches using VSD. First, we observe training instability across different architectures, primarily due to the reliance on a fixed guidance scale. Second, we identify a broader issue affecting all one-step generators: the lack of negative prompt guidance, which limits the ability to exclude specific elements from generated images—a feature commonly supported in multi-step models [15, 8, 20]. These issues limit the flexibility and control of one-step models, which reduces their practical effectiveness despite their faster generation times.
Prior research [54] has attempted to address the stability issues of the VSD framework, but these solutions come with their own limitations. Techniques like increased teacher update frequency [54] and adversarial loss terms [42, 19] have been explored to stabilize the training process. However, these approaches can add a significant computational burden, partly due to the additional proposed technique itself and partly due to the introduction of image supervision. Since we’re focusing on building an image-free distillation, given these limitations, there is a pressing need to conduct a deeper analysis of the VSD framework to understand its challenges and develop more efficient, stable solutions.
Alongside these stability limitations, as previously mentioned, one-step diffusion models currently lack support for negative prompting, a feature widely used in multi-step diffusion models to enhance image quality and filter out unwanted features. In multi-step models, negative prompting techniques are particularly useful for refining alignment with text prompts in text-to-2D generation, and they also help address issues like the Janus problem in text-to-3D generation [8, 20, 1] —where models produce mirrored or duplicated features like faces. One common way to integrate negative prompts into multi-step diffusion model sampling is using Classifier-free Guidance (CFG) [15]. Given two noise predictions from the positive prompt and negative prompt, CFG subtracts negative noise prediction from the positive noise prediction iteratively throughout the denoising process, inherently steering the final output away from the negative prompt. However, one-step models do not have the iterative refinement process required by CFG, making it difficult to apply negative prompting in these frameworks. This limitation reduces the flexibility and control of one-step models. To address this, a novel approach is needed to incorporate the negative prompt signal directly within the one-step generation process, enabling these models to effectively avoid unwanted features without iterative sampling. We address these challenges by introducing a robust framework called SNOOPI for the one-step model, illustrated in Fig. 1, that enhances stability and control during both training and inference. (1) We propose Proper Guidance - SwiftBrush (PG-SB), a dynamic guidance approach that varies the guidance scale of teacher models during training to stabilize VSD. This flexible CFG improves the adaptability of image-free distillation across diverse diffusion backbones without adding computational or memory overhead. (2) We introduce a method called Negative-Away Steer Attention (NASA) used in inference, which incorporates negative prompt guidance into one-step models by adjusting cross-attention layers, effectively reducing unwanted features in generated images. By steering attention within intermediate feature space, NASA offers image control capabilities that were previously achievable only in multi-step diffusion models.
Our contributions are threefold:
  • We propose Proper Guidance-SwiftBrush (PG-SB), applying a varied guidance scale to teacher model outputs during training, to enhance stability across different diffusion backbones without additional computational costs.
  • We introduce Negative-Away Steer Attention (NASA), the first method to integrate negative prompt guidance into one-step diffusion models, effectively suppressing unwanted visual attributes.
  • Extensive experimental results on established benchmarks demonstrate that our method outperforms strong baselines, validating its stability and effectiveness.
2 Related Work
2.1 Text-to-image Diffusion Distillation
Recent advances in text-to-image diffusion models have focused on accelerating generation through several key methodological streams: conventional distillation, trajectory optimization, and distribution matching.
Conventional Distillation. Early work in model distillation explored two main directions. Progressive distillation methods [40, 32] iteratively train a sequence of student models, each halving the required sampling steps while preserving generation quality. While effective, this approach necessitates multiple training phases, increasing overall computational cost. In parallel, direct distillation techniques [28, 12] generate synthetic noise-image pairs using a teacher model to train students via reconstruction loss, offering a more straightforward training process but often struggling with quality retention. Recent work has augmented these methods with adversarial training [19, 42], achieving better image quality at the cost of increased training complexity.
Trajectory Optimization. Trajectory-straightening methods [25, 26] simplify the complex denoising trajectories into near-linear paths, enabling faster sampling without compromising quality. Meanwhile, consistency-based frameworks [46, 21, 44] use self-consistency properties to train models to align outputs across multiple points in the ODE-based denoising trajectory.
Distribution Matching. More recent work [33, 7, 55, 54] has explored distribution matching techniques. These methods directly align the output distributions of teacher and student models, particularly through Variational Score Distillation (VSD). These techniques can be broadly categorized into two approaches: image-dependent training and image-free training.
Image-dependent methods, which rely on real images for adversarial loss [54, 41] or synthetic ones generated by the teacher models for reconstruction loss [55]. Despite their effectiveness in preserving image quality, these approaches face significant practical limitations: they require substantial storage capacity, consume high memory during training, demand extensive computational resources, and risk mode collapse due to intensive image supervision.
In contrast, image-free approaches [33, 7] have demonstrated promising results by relying solely on prompt text for training, offering a significant reduction in resource requirements while maintaining competitive performance. However, image-free techniques often encounter training instability, particularly within VSD-based frameworks. Existing solutions to stabilize training [55] introduce significant memory and computational overhead, limiting their practical application. Our work addresses this gap by introducing a simple yet effective method that enhances training stability without additional computational burden, making image-free variational score distillation more accessible for real-world applications.
2.2 Sampling Guidance and Negative Prompting
Classifier-free guidance (CFG) [15] and negative prompting are key techniques for content control in image generation, from text-to-2D to text-to-3D synthesis. In multi-step diffusion models, these methods enhance desired image features and suppress undesired ones, improving visual fidelity and reducing artifacts like the Janus problem in 3D tasks [8, 20, 1]. Score-based approaches like NFSD [20] employ negative prompting to steer generated samples toward real-image distributions, while SDS-Bridge [31] uses negative prompts to model source distributions more accurately. Despite their success, negative prompting has yet to be integrated into one-step or few-step diffusion models. To bridge this gap, we introduce Negative-Away Steer Attention (NASA), the first method to enable negative prompting in one-step and few-step diffusion models, enhancing control in faster generation settings.
3 Background
Diffusion Models [16, 45, 38] produce high-quality images by progressively denoising inputs. This technique involves a forward process that gradually adds noise to data and a reverse process that reconstructs the original data by removing noise. While the original diffusion models act on image space [16, 45], LDM [38] processes on latent space produced by a pretrained VAE for more efficient computation. In particular, starting with a data point $\mathbf{x}_{0}$ sampled from an unknown distribution $q(\mathbf{x}_{0})$, the forward process gradually diffuses $\mathbf{x}_{0}$ into a standard Gaussian noise $\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ through $T$ consecutive timesteps, where $\mathbf{I}$ is the identity matrix. At each time-step $t$, a noisier version of $\mathbf{x}_{0}$, denoted as $\mathbf{x}_{t}$, is drawn from $q_{t}(\mathbf{x}_{t}|\mathbf{x}_{0})=\mathcal{N}(\alpha_{t}\mathbf{x}_{0},\sigma^{2}_{t}\mathbf{I})$ using a standard Gaussian noise $\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$:
$\mathbf{x}_{t}=\alpha_{t}\mathbf{x}_{0}+\sigma_{t}\boldsymbol{\epsilon},\quad\forall t\in\{0,\dots,T\},$ (1)
where $\{(\alpha_{t},\sigma_{t})\}_{t=1}^{T}$ defines the noise schedule, with boundary conditions $(\alpha_{0},\sigma_{0})=(1,0)$ for the clean sample and $(\alpha_{T},\sigma_{T})=(0,1)$ for pure noise.
Conversely, the reverse process aims to gradually reconstruct the original data by denoising the input over $T$ steps, starting from an initial random Gaussian noise $\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. The model is trained by minimizing the difference between the noise estimated by the model, $\boldsymbol{\epsilon}_{\phi}$, parameterized by $\phi$, and the actual noise in Eq. 1:
$\min_{\phi}\mathbb{E}_{t\sim\mathcal{U}(0,T),\boldsymbol{\epsilon}\sim\mathcal{N}(0,I)}\|\boldsymbol{\epsilon}_{\phi}(\mathbf{x}_{t},t)-\boldsymbol{\epsilon}\|_{2}^{2}.$ (2)
Text-to-Image Diffusion Models guide the sampling process using an additional text prompt $\mathbf{y}$ to produce outputs that are both photorealistic and aligned with the provided textual descriptions. The training objective, slightly modified from the unconditional loss in Eq. 2, is defined as:
$\min_{\phi}\mathbb{E}_{t\sim\mathcal{U}(0,T),\boldsymbol{\epsilon}\sim\mathcal{N}(0,I),\mathbf{y}}\|\boldsymbol{\epsilon}_{\phi}(\mathbf{x}_{t},t,\mathbf{y})-\boldsymbol{\epsilon}\|_{2}^{2}.$ (3)
Classifier-free Guidance (CFG) [15] is an inference technique designed to enhance the quality of generated images by blending the predictions from both a conditional and an unconditional model. At each sampling step, CFG adjusts the denoiser’s output using a control parameter $\kappa>1$, allowing for controlled guidance that aligns more closely with the desired conditions:
$\hat{\boldsymbol{\epsilon}}_{\phi}(\mathbf{x}_{t},t,{\bf y},\kappa)=(1-\kappa)\,\boldsymbol{\epsilon}_{\phi}(\mathbf{x}_{t},t)+\kappa\,\boldsymbol{\epsilon}_{\phi}(\mathbf{x}_{t},t,{\bf y}).$ (4)
Negative Prompts provide enhanced control by suppressing unwanted features in the generated content. Instead of producing an unconditional output, the model generates an output conditioned on the negative prompt $\mathbf{y}^{-}$, as follows:
$\hat{\boldsymbol{\epsilon}}_{\phi}(\mathbf{x}_{t},t,\mathbf{y},\mathbf{y}^{-},\kappa)=(1-\kappa)\,\boldsymbol{\epsilon}_{\phi}(\mathbf{x}_{t},t,\mathbf{y}^{-})+\kappa\,\boldsymbol{\epsilon}_{\phi}(\mathbf{x}_{t},t,\mathbf{y}).$ (5)
Variational Score Distillation (VSD) is a powerful framework that utilizes pretrained diffusion-based text-to-image models to enhance text-based generation across both 3D and 2D domains. Originally proposed for text-to-3D tasks [48], VSD leverages diffusion-based score matching to align Neural Radiance Fields (NeRFs) with text prompt, enabling the generation of intricate 3D objects [48]. This approach extends effectively to 2D image synthesis as well, where VSD enables rapid, high-quality text-to-image generation in a single step with models such as SwiftBrush and DMD [33, 7, 55, 54], sidestepping the computational burden of multi-step diffusion methods.
The central aim of VSD is to ensures that the renderings of a differentiable generator align with the probability density of plausible images as guided by the 2D diffusion model. To accomplish this, VSD employs a two-teacher approach that uses a fixed pretrained diffusion model $\boldsymbol{\epsilon}_{\psi}$ and an adaptive LoRA-based teacher $\boldsymbol{\epsilon}_{\phi}$. While training, the student model $\boldsymbol{f}_{\theta}$ produces 2D images $\hat{\mathbf{x}}_{0}=\boldsymbol{f}_{\theta}(\mathbf{z},\mathbf{y})$ using an input noise $\mathbf{z}\sim\mathcal{N}(0,I)$ and a text prompt $\mathbf{y}$. Noisy images $\hat{\mathbf{x}}_{t}=\alpha_{t}\hat{\mathbf{x}}_{0}+\sigma_{t}\boldsymbol{\epsilon}$ is then fed into both teacher models. The LoRA teacher model $\boldsymbol{\epsilon}_{\phi}$ aligns with the student distribution by minimizing a denoising L2 loss on single-step samples. This arrangement supports robust and adaptive guidance suitable for a variety of generative architectures. The gradient of the learning objective with respect to the student model parameters $\theta$ is defined as:
$\begin{split}\nabla_{\theta}\mathcal{L}_{\text{VSD}}=\mathbb{E}_{t,\mathbf{y},\mathbf{z},\kappa=C}&\bigg{[}\,w(t)\left(\hat{\boldsymbol{\epsilon}}_{\psi}(\hat{\mathbf{x}}_{t},t,\mathbf{y},\kappa)\right.\\ &\left.-\hat{\boldsymbol{\epsilon}}_{\phi}(\hat{\mathbf{x}}_{t},t,\mathbf{y},\kappa\right)\frac{\partial\boldsymbol{f}_{\theta}(\mathbf{z},\mathbf{y})}{\partial\theta}\bigg{]},\end{split}$ (6)
where $C$ is a constant and $w(t)$ is the time-dependent weight that adjusts the contribution of each timestep, aligning the student’s outputs with the diffusion model’s predictions. Through alternating updates of the student and LoRA teacher, VSD enables efficient, high-fidelity generation across both text-to-image and text-to-3D tasks.
4 Proper Guidance - SwiftBrush
Figure 2: FID 2K progression throughout training. The default baseline is SwiftBrush for both SDv1.5 and PixArt-$\alpha$. DMD2Trick indicates the Two Time-scale Update Rule technique [54], which increases the update frequency of the LoRA teacher to 5 times after each student update. These baselines demonstrate notable instability, leading to variability in image quality metrics such as FID. In contrast, our approach, which employs CFG with randomly selected values from a uniform distribution, results in a more stable training process and generates higher-quality samples.
4.1 Motivation
While the VSD framework offers a solid foundation for model distillation, its training process often experiences instability. For example, while successfully working on Stable Diffusion 2.1 (SDv2.1) [38], our re-implementation of SwiftBrush and SwiftBrushv2 (SB) [33, 7] fail in distilling other model backbones such as Stable Diffusion 1.5 (SDv1.5) [38] and PixArt-$\alpha$ [6] when using a fixed CFG scale in the VSD framework. Their unstable results are substantiated by quantitative evaluation shown in Fig. 2.
Figure 3: Qualitative comparison of different methods for integrating negative prompts into the generation process using a one-step generator. As shown, simply applying CFG to a one-step generator is equivalent to blending (i.e., adding) two images together, resulting in an unusable output. Alternatively, negating the negative embedding from the positive embedding shows minimal impact on the final image. In contrast, NASA is the first method to successfully steer the generation away from negative attributes in a one-step generator, while also producing high-quality results.
To address this instability, previous studies have proposed techniques such as using image datasets as regularizers, increasing the LoRA teacher update frequency, and adding adversarial loss terms [54, 55]. However, these methods have two main drawbacks: (1) they rely on image supervision, and (2) they significantly increase memory usage and training time. In pursuit of image-free distillation, we instead investigate the existing SB framework to identify the root cause of the instability.
We begin by examining the impact of the CFG scale in multi-step diffusion models, specifically SDv2.1. Our analysis of the MS COCO 2014 benchmark [24], with results shown in Tab. 1, reveals that model performance shifts considerably as the guidance scale changes. Typically, users set the guidance scale based on their specific goal, whether to prioritize image fidelity or text alignment. However, this approach is less desirable in training, as we want the model to be versatile across different scenarios. Given this, we seek to make the teacher model adaptable across various scenarios without compromising overall performance. As shown in Tab. 1, varying the guidance scale randomly allows us to achieve competitive results across all metrics while maintaining flexibility across different guidance levels.
$\kappa$ FID $\downarrow$ CLIP $\uparrow$ Precision $\uparrow$ Recall $\uparrow$
2 10.77 0.31 0.54 0.53
3 9.60 0.32 0.59 0.50
4 10.35 0.32 0.60 0.47
5 11.53 0.33 0.61 0.44
$\mathcal{U}(2,5)$ 9.86 0.32 0.60 0.48
Table 1: Results on the zero-shot MS COCO-2014 30K benchmark [24] of SDv2.1 [38] when using PNDM Scheduler with different CFG scale $\kappa$.
Teacher LoRA Teacher FID $\downarrow$
10.72
10.81
9.51
9.03
Table 2: Ablation studies on Random CFG in VSD loss with SDv2.1 backbone on 2K prompts from MS COCO-2014 [24].
4.2 Method
Based on our findings, we aim to broaden the guidance distribution provided by the teacher in the SB distillation framework by introducing Proper Guidance - SwiftBrush (PG-SB) which applies a randomized guidance scale during teacher inference. Specifically, rather than fixing $\kappa=C$, we sample its value from a uniform distribution $\boldsymbol{\mathcal{U}(\kappa_{\text{min}},\kappa_{\text{max}})}$ where $\kappa_{\text{min}}$ and $\kappa_{\text{max}}$ are predefined constants. In the SB distillation pipeline, however, CFG is applied in two places: once to the frozen teacher and once to the LoRA teacher. As shown in Tab. 2, applying this technique only to the LoRA teacher improves model performance. However, applying it solely to the frozen teacher destabilizes the distillation process and eventually harms the student performance. The gradient of our PG-SB’s VSD loss is defined by:
$\begin{split}\nabla_{\theta}\mathcal{L}_{\text{VSD}}^{\text{PG-SB}}=&\mathbb{E}_{t,\mathbf{y},\mathbf{z},\boldsymbol{\kappa\sim\mathcal{U}(\kappa_{\text{min}},\kappa_{\text{max}})}}\bigg{[}\,w(t)\left(\hat{\boldsymbol{\epsilon}}_{\psi}(\hat{\mathbf{x}}_{t},t,\mathbf{y},\kappa)\right.\\ &\left.-\hat{\boldsymbol{\epsilon}}_{\phi}(\hat{\mathbf{x}}_{t},t,\mathbf{y},\kappa\right)\frac{\partial\boldsymbol{f}_{\theta}(\mathbf{z},\mathbf{y})}{\partial\theta}\bigg{]}.\end{split}$ (7)
Figure 4: Left: An overview of the Negative-Away Steer Attention (NASA) pipeline. Positive (green) and negative (red) prompts are fed into a text encoder to generate positive and negative text features. The NASA module then processes these features, which adjusts the one-step diffusion model to steer the output image away from the negative features, refining it based on the positive features. Right: The details of the NASA module. It processes queries ($\mathbf{Q}_{l}$) in layer $l$, note we will omit the subscript $l$ in subsequent notations to improve readability, with positive ($\mathbf{V}^{+},\mathbf{K}^{+}$) and negative ($\mathbf{V}^{-},\mathbf{K}^{-}$) key-value pairs to create positive ($\mathbf{Z}^{+}$) and negative ($\mathbf{Z}^{-}$) attention outputs. The final output ($\mathbf{Z}^{\text{NASA}}$) is calculated by subtracting the weighted negative features ($\mathbf{Z}^{-}$) from the positive features ($\mathbf{Z}^{+}$).
We hypothesize that broadening the output distribution of the LoRA teacher enables it to better align with the student’s output, mainly since the student may not initially perform well at one function evaluation or one NFE generation. Interestingly, once the LoRA teacher’s output distribution is expanded, the student can also adapt to the broader range of outputs from the frozen teacher when using a randomized guidance scale, achieving the best results when varied guidance scales are applied to both teachers.
5 Negative-Away Steer Attention
5.1 Motivation
Negative prompting is an essential technique for improving control in text-to-image and text-to-3D synthesis. This method helps exclude unwanted elements in generated images and mitigate the Janus problem in 3D generation [20, 31, 1]. However, no existing approach supports negative prompts in single-step or few-step diffusion models.
A widely method for implementing negative prompting relies on the CFG mechanism. However, applying CFG directly to one-step diffusion models presents significant challenges. CFG was designed for multi-step models, where the output is iteratively refined, allowing for gradual adjustments using both negative and unconditional outputs. In contrast, one-step models generate images in a single pass without the iterative denoising process, making them inherently less flexible. Directly applying CFG to these models leads to an undesirable blending of images, resulting in unnatural artifacts, as shown in Fig. 3. This highlights a crucial limitation: the CFG mechanism, though effective in multi-step models, is not directly translatable to one-step models without significant modifications.
An alternative approach manipulates text embeddings by subtracting the negative embedding from the positive one to steer the model away from undesired attributes. However, as shown in Fig. 3, this method has minimal effect on the generated images, highlighting the need for a more robust solution to negative prompting in one-step models.
5.2 Method
Instead of applying CFG in the input space (text embedding) or output space (synthesized image) as in multi-step diffusion, we propose moving CFG to the intermediate feature space of the diffusion model. We hypothesize that cross-attention layer features are particularly well-suited for this approach. Specifically, cross-attention layers in these models capture semantic connections between image regions and text features [13], making them ideal for controlling feature emphasis. By modulating these cross-attention maps, we can filter unwanted features by subtracting attention responses to negative prompts from those to positive prompts. This filtering occurs in the feature space, meaning that undesired elements are excluded before the image is generated, reducing the chance of blending artifacts.
Given the latent’s features at current layer $l$, $\mathbf{Z}_{l}$, and positive text features $\mathbf{c}_{p}$ extracted from CLIP text encoder, we define the query, key, and value matrices as $\mathbf{Q}_{l}=\mathbf{Z}_{l}\mathbf{W}_{q,l}$, $\mathbf{K}_{l}^{+}=\mathbf{c}_{p}\mathbf{W}_{k,l}$, and $\mathbf{V}_{l}^{+}=\mathbf{c}_{p}\mathbf{W}_{v,l}$, where $\mathbf{W}_{q,l}$, $\mathbf{W}_{k,l}$, and $\mathbf{W}_{v,l}$ are projection matrices at current layer $l$. For negative text features $\mathbf{c}_{n}$, we use the same projection matrices, defining the key and value matrices as $\mathbf{K}_{l}^{-}=\mathbf{c}_{n}\mathbf{W}_{k,l}$ and $\mathbf{V}_{l}^{-}=\mathbf{c}_{n}\mathbf{W}_{v,l}$.
The NASA mechanism is formulated as follows:
$\mathbf{Z}_{l}^{+}$ $=\text{Softmax}\left(\frac{\mathbf{Q}_{l}(\mathbf{K}_{l}^{+})^{\top}}{\sqrt{d}}\right)\mathbf{V}_{l}^{+},$ (8)
$\mathbf{Z}_{l}^{-}$ $=\text{Softmax}\left(\frac{\mathbf{Q}_{l}(\mathbf{K}_{l}^{-})^{\top}}{\sqrt{d}}\right)\mathbf{V}_{l}^{-},$ (9)
$\mathbf{Z}^{\text{NASA}}_{l}=\mathbf{Z}_{l}^{+}-\alpha\cdot\mathbf{Z}_{l}^{-},$ (10)
where $\alpha$ is the scale factor to control the degree of negative feature removal. The overall pipeline of the NASA mechanism is illustrated in Fig. 4.
6 Experiments
6.1 Experimental Setup
Metrics and benchmark. In diffusion model evaluation, Fréchet Inception Distance (FID) [14] on the zero-shot MS-COCO2014 benchmark has traditionally been the standard metric. However, recent studies [37, 5, 54] suggest that FID does not always correlate with actual image quality and adopt human-focused alignment metrics like Human Preference Score v2 (HPSv2) [49] for benchmarking. Our experiments also follow this conclusion, so we choose HPSv2 as the primary metric for evaluating the models. Additionally, we use CLIP scores to assess text alignment and precision/recall [22] to evaluate general image quality and mode coverage. For completeness, we still report FID results. Except for HPSv2, all metrics are computed following the protocols in [18, 7]: generating 30K images using MS-COCO2014 prompts, resizing them to $256\times 256$, and comparing with center-cropped images from the full dataset.
Implementation details. We base our method on SBv2 [7] with our proposed modifications on CFG scale $\kappa$. We use total of 3.3M prompts from JourneyDB [47] and a subset of LAION [43]. All our training is conducted on four NVIDIA A100 40GB GPUs with the total batch size of 64. The learning rates are set to $1e^{-6}$ for the student model and $1e^{-3}$ for the LoRA teacher, using the AdamW optimizer [27]. We utilize three backbones SDv1.5 [38], SDv2.1 [38] and PixArt-$\alpha$ [6] as the frozen teachers and LoRA teachers are initialized with rank $r=64$ and scaling $\gamma=128$. The teacher models’ CFG scales are set as follows: $\kappa_{\text{min}}=0.5,\kappa_{\text{max}}=4$ for SDv1.5 and SDv2.1 and $\kappa_{\text{min}}=0.5,\kappa_{\text{max}}=3$ for PixArt-$\alpha$. Further details can be found in the Appendix.
6.2 Evaluation of PG-SB
Tab. 4 and Tab. 5 present a thorough quantitative comparison between our approach and prior distilled text-to-image diffusion models. The results consistently show the superiority of our model in HPSv2 scores, with the highest average score across all tested backbones. This reinforces the robustness of our method, as discussed in Sec. 4.2 and demonstrated in Fig. 2, while also revealing a key insight. By gaining a deep understanding of the underlying framework, we made a simple yet impactful change with minimal overhead, leading to significant performance improvements. This suggests that integrating complex loss functions may not always be necessary for achieving state-of-the-art results. Remarkably, our method achieves these outcomes without requiring image-level supervision, restoring SBv2 as an image-free distillation approach.
NASA SDXL-LCM SDXL-DMD2 SDXL-DMD2 PG-SB
4 steps 4 steps 1 step 1 step
43% 27% 25% 38%
97% 100% 100% 92%
Table 3: Comparison of success rate of unwanted feature removal in generated images between models with and without NASA.
Figure 5: Effect of different scale values (0.0 to 1.0) in NASA with SDXL-DMD2 model, illustrating the progressive influence on visual details and composition.
Figure 6: Qualitative results of NASA. Each row displays images generated with a positive (green) and negative (red) prompt pair. Models using NASA effectively exclude unwanted features, aligning outputs with desired traits.
Method Anime Photo Concept Art Paintings Average
Stable Diffusion 1.5-based backbone
SDv1.5 [38] 26.51 27.19 26.06 26.12 26.47
InstaFlow-0.9B [26] 26.10 26.62 25.92 25.95 26.15
DMD2 [54] 26.39 27.00 25.80 25.83 26.26
PG-SB 27.18 27.58 26.69 26.62 27.02
PG-SB + NASA 27.19 27.59 26.71 26.63 27.03
Stable Diffusion 2.1-based backbone
SDv2.1 [38] 27.48 26.89 26.86 27.46 27.17
SB [33] 26.91 27.21 26.32 26.37 26.70
SBv2 [7] 27.25 27.62 26.86 26.77 27.13
PG-SB 27.56 27.84 26.97 27.03 27.35
PG-SB + NASA 27.71 27.99 27.14 27.27 27.53
PixArt-$\alpha$-based backbone
PixArt-$\alpha$ [6] 29.62 29.17 28.79 28.69 29.07
YOSO [30] 28.79 28.09 28.57 28.55 28.50
DMD [55] 29.31 28.67 28.46 28.41 28.71
PG-SB 32.19 29.09 30.39 29.69 30.34
PG-SB + NASA 32.56 29.55 31.24 30.96 31.08
Table 4: HPSv2 comparisons between our method and previous works. denotes reported numbers, denotes our rerun based on the publicly available model checkpoints.
6.3 Evaluation of NASA
To assess the effectiveness of NASA, we conduct experiments with our model and other one-step and few-step models, such as LCM [29] and DMD2 [54]. We design a small set of positive and negative prompts and generate 100 images per model, both with and without negative prompts (further details are provided in the Appendix). We then calculate the percentage of images that successfully exclude the feature specified by the negative prompt. For example, given the positive prompt “A photo of a person” and the negative prompt “male”, we measure the percentage of generated images that do not depict a male person. The average results across all prompt sets are summarized in Tab. 3. Our findings show that NASA is compatible with both one-step and few-step models, enabling precise control over unwanted features in generated images. Note that the difference in accuracy between PG-SB and other models is primarily attributed to the disparity in backbone size. The SDXL-based architecture is well-known for its superior text alignment, which stems from its larger generator backbone and the use of an ensemble of text encoders. Additionally, Fig. 5 and Fig. 6 present qualitative results of our method. Images generated with a conditional negative prompt retain key semantic features such as pose and background, while filtering out unwanted attributes specified by the negative prompt. Furthermore, when using negative prompts like “worst quality, low quality, ugly, duplicate, out of frame, deformed, blurry, bad anatomy, watermark”, our method improves image quality, as shown in Tab. 4 and Tab. 5. Notably, integrating NASA with PG-SB on the PixArt-$\alpha$-based backbone achieves a record-breaking HPSv2 score of 31.08, surpassing all other one-step diffusion models baselines.
Method #Params NFEs FID $\downarrow$ CLIP $\uparrow$ Precision $\uparrow$ Recall $\uparrow$ Image-free?
Stable Diffusion 1.5-based backbone
SDv1.5 ($\textit{cfg}=3$) [38] 0.9B 25 8.78 0.30 0.59 0.53
UFOGen [51] 0.9B 1 12.78 - - -
InstaFlow-0.9B [26] 0.9B 1 13.33 0.30 0.53 0.45
DMD [55] 0.9B 1 11.49 0.32 - -
DMD2 [54] 0.9B 1 8.29 0.30 0.59 0.52
PG-SB 0.9B 1 10.08 0.31 0.57 0.47
PG-SB + NASA 0.9B 1 9.94 0.31 0.57 0.48
Stable Diffusion 2.1-based backbone
SDv2.1 ($\textit{cfg}=3$) [38] 0.9B 25 9.60 0.32 0.59 0.50
SB [33] 0.9B 1 15.46 0.30 0.47 0.46
SBv2 [7] 0.9B 1 8.14 0.32 0.57 0.52
PG-SB 0.9B 1 8.91 0.34 0.57 0.49
PG-SB + NASA 0.9B 1 8.83 0.34 0.57 0.50
PixArt-$\alpha$-based backbone
PixArt-$\alpha$ ($\textit{cfg}=4.5$) [6] 0.6B 20 26.85 0.32 0.52 0.23
YOSO [30] 0.6B 1 26.04 0.30 0.46 0.30
DMD [55] 0.6B 1 30.22 0.32 0.55 0.18
PG-SB 0.6B 1 22.03 0.33 0.52 0.24
PG-SB + NASA 0.6B 1 21.98 0.33 0.52 0.25
Table 5: Quantitative comparisons between our method and others on zero-shot MS COCO-2014 benchmark. For multi-step SD models, we report each with the CFG scale that returns the best FID. denotes reported numbers, denotes our rerun based on the publicly available model checkpoints. ‘-’ denotes unreported results.
7 Discussion and Conclusion
Limitations: Despite its promising results, PG-SB currently lacks support for few-step models, limiting its applicability to downstream tasks. Additionally, while NASA proves effective, achieving high-quality images depends on selecting an appropriate scale for removing negative features, though the parameter does not require extensive tuning. Furthermore, NASA is designed specifically for architectures with cross-attention layers, leaving support for other architectures, such as MM-DiT [9], as future work.
Conclusion: This paper introduces SNOOPI, a framework that improves stability and control in one-step diffusion models for both training and inference. PG-SB dynamically adjusts the guidance scale to stabilize VSD across various model backbones without increasing computational cost. Additionally, NASA is the first approach to incorporate negative prompt guidance in one-step diffusion models, enabling the effective reduction of unwanted features in generated images through cross-attention adjustments. Experimental results demonstrate that SNOOPI outperforms strong baselines, showcasing its stability and effectiveness.
Supplementary Material
8 Implementation Details
PG-SB: Following SBv2 [7], we also use the clamped CLIP loss with a margin of $\tau=0.37$, starting with a weight of 0.1 and gradually reducing to zero. Tab. 6 provides the additional hyperparameters for training PG-SB.
Hyperparameter SDv1.5 SDv2.1 PixArt-$\alpha$
Dataset JDB + LAION JDB + LAION JDB
Batch size 64 64 64
Training iterations 60k 40k 60k
Mixed-Precision (BF16) Yes Yes Yes
$(\kappa_{\text{min}},\kappa_{\text{max})}$ (0.5, 4) (0.5, 4) (0.5, 3)
Clip weight 0.1 0.1 0.1
$\tau$ 0.37 0.37 0.37
$lr$ of student $1e^{-6}$ $1e^{-6}$ $1e^{-6}$
$lr$ of LoRA teacher $1e^{-3}$ $1e^{-3}$ $1e^{-3}$
LoRA rank $r$ 64 64 64
LoRA scaling $\gamma$ 128 128 128
Table 6: Hyperparameters used for training PG-SB.
Positive prompt Negative prompt SDXL-LCM SDXL-DMD2 SDXL-DMD2 PG-SB
4 steps 4 steps 1 step 1 step
“A photo of a person” “male” 52% $\rightarrow$ 99% 26% $\rightarrow$ 100% 29% $\rightarrow$ 100% 42% $\rightarrow$ 81%
“female” 48% $\rightarrow$ 100% 74% $\rightarrow$ 100% 71% $\rightarrow$ 100% 58% $\rightarrow$ 88%
“A photo of a person” “young” 55% $\rightarrow$ 100% 44% $\rightarrow$ 100% 7% $\rightarrow$ 100% 39% $\rightarrow$ 96%
“old” 45% $\rightarrow$ 92% 56% $\rightarrow$ 100% 93% $\rightarrow$ 100% 61% $\rightarrow$ 79%
“A photo of a pet” “cat” 36% $\rightarrow$ 100% 12% $\rightarrow$ 100% 60% $\rightarrow$ 100% 32% $\rightarrow$ 99%
“dog” 64% $\rightarrow$ 100% 88% $\rightarrow$ 100% 40% $\rightarrow$ 100% 68% $\rightarrow$ 99%
Table 7: Comparison of success rates of unwanted feature removal in generated images before and after applying the NASA model.
Method Anime Photo Concept Art Paintings Average
Stable Diffusion 1.5-based backbone
InstaFlow-0.9B [26] 26.10 26.62 25.92 25.95 26.15
InstaFlow-0.9B + NASA ($\alpha=0.19$) 26.24 26.74 26.04 26.00 26.26 (+0.11)
DMD2 [54] 26.39 27.00 25.80 25.83 26.26
DMD2 + NASA ($\alpha=0.04$) 26.41 27.02 25.80 25.83 26.27 (+0.01)
PG-SB 27.18 27.58 26.69 26.62 27.02
PG-SB + NASA ($\alpha=0.04$) 27.19 27.59 26.71 26.63 27.03 (+0.01)
Stable Diffusion 2.1-based backbone
SBv2 [7] 27.25 27.62 26.86 26.77 27.13
SBv2 + NASA ($\alpha=0.30$) 27.45 27.85 26.93 27.09 27.33 (+0.20)
PG-SB 27.56 27.84 26.97 27.03 27.35
PG-SB + NASA ($\alpha=0.30$) 27.71 27.99 27.14 27.27 27.53 (+0.18)
PixArt-$\alpha$-based backbone
YOSO [30] 28.79 28.09 28.57 28.55 28.50
YOSO + NASA ($\alpha=0.20$) 28.80 28.10 28.62 28.57 28.52 (+0.02)
DMD [55] 29.31 28.67 28.46 28.41 28.71
DMD + NASA ($\alpha=0.45$) 29.34 28.71 28.50 28.52 28.77 (+0.06)
PG-SB 32.19 29.09 30.39 29.69 30.34
PG-SB + NASA ($\alpha=0.80$) 32.56 29.55 31.24 30.96 31.08 (+0.74)
Table 8: HPSv2 comparisons between our method and previous works. denotes reported numbers, denotes our rerun based on the publicly available model checkpoints.
NASA: We provide the PyTorch pseudo-code for the NASA Attention algorithm, as outlined in Sec. 5. The implementation is straightforward and is provided in Algorithm 1.
Algorithm 1 NASA Attention.
class NASA_AttnProcessor(nn.Module):
def __init__(self, scale=1.0):
super().__init__()
self.scale = scale
def __call__(
self, attn, z, emb, neg_emb, attn_mask, neg_attn_mask, temb, scale=1., *args, **kwargs,
# Input preparation
residual = z
if attn.spatial_norm is not None:
z = attn.spatial_norm(z, temb)
input_ndim = z.ndim
if input_ndim == 4:
bsz, channel, height, width = z.shape
z = z.view(bsz, channel, height * width).transpose(1, 2)
bsz, sequence_length, _ = z.shape if emb is None else emb.shape
_, neg_sequence_length, _ = z.shape if neg_emb is None else neg_emb.shape
if attn_mask is not None:
attn_mask = attn.prepare_attention_mask(attn_mask, sequence_length, bsz)
attn_mask = attn_mask.view(bsz, attn.heads, -1, attn_mask.shape[-1])
if neg_attn_mask is not None:
neg_attn_mask = attn.prepare_attention_mask(neg_attn_mask, neg_sequence_length, bsz)
neg_attn_mask = neg_attn_mask.view(bsz, attn.heads, -1, neg_attn_mask.shape[-1])
if attn.group_norm is not None:
z = attn.group_norm(z.transpose(1, 2)).transpose(1, 2)
query = attn.to_q(z)
if emb is None:
emb = z
else:
if attn.norm_cross:
emb = attn.norm_emb(emb)
neg_emb = attn.norm_emb(neg_emb)
# Compute cross-attention for positive embedding
key = attn.to_k(emb)
value = attn.to_v(emb)
inner_dim = key.shape[-1]
head_dim = inner_dim // attn.heads
query = query.view(bsz, -1, attn.heads, head_dim).transpose(1, 2)
key = key.view(bsz, -1, attn.heads, head_dim).transpose(1, 2)
value = value.view(bsz, -1, attn.heads, head_dim).transpose(1, 2)
if attn.norm_q is not None:
query = attn.norm_q(query)
if attn.norm_k is not None:
key = attn.norm_k(key)
z = F.scaled_dot_product_attention(
query, key, value, attn_mask=attn_mask, dropout_p=0.0, is_causal=False
z = z.transpose(1, 2).reshape(bsz, -1, attn.heads * head_dim)
z = z.to(query.dtype)
# Compute cross-attention for negative embedding
neg_key = attn.to_k(neg_emb)
neg_value = attn.to_v(neg_emb)
neg_key = neg_key.view(bsz, -1, attn.heads, head_dim).transpose(1, 2)
neg_value = neg_value.view(bsz, -1, attn.heads, head_dim).transpose(1, 2)
neg_z = F.scaled_dot_product_attention(
query, neg_key, neg_value, attn_mask=neg_attn_mask, dropout_p=0.0, is_causal=False
neg_z = neg_z.transpose(1, 2).reshape(bsz, -1, attn.heads * head_dim)
neg_z = neg_z.to(query.dtype)
# Equation 10 in main paper
z_nasa = z - self.scale * neg_z
z_nasa = attn.to_out[0](z_nasa)
z_nasa = attn.to_out[1](z_nasa)
if input_ndim == 4:
z_nasa= z_nasa.transpose(-1, -2).reshape(bsz, channel, height, width)
if attn.residual_connection:
z_nasa = z_nasa + residual
z_nasa = z_nasa / attn.rescale_output_factor
return z_nasa
9 More Analysis on NASA
9.1 Experimental Details in Tab. 3
For clarification, in Tab. 3, we evaluate the effectiveness of NASA across four models: SDXL-LCM, SDXL-DMD2 (with both 1-step and 4-step settings), and our PG-SB model using the SDv2.1 backbone. For the “without NASA” case, all models use a scale of $\alpha=0$, while for the “with NASA” case, a scale of $\alpha=0.5$ is applied. The evaluation involves 6 positive-negative prompt pairs. For each model and prompt pair, 100 images are generated. Human evaluations are then used to determine the percentage of images that successfully exclude the feature specified by the negative prompt. The detailed results from Tab. 3 are provided in Tab. 7.
9.2 More Quantitative Results
Tab. 8 presents additional quantitative results when applying NASA to various models. Using a negative prompt such as “worst quality, low quality, ugly, duplicate, out of frame, deformed, blurry, bad anatomy, watermark,” our NASA method demonstrates a slight improvement in the HPSv2 scores across all models.
10 Additional Qualitative Results of PG-SB
Fig. 7, Fig. 8, and Fig. 9 present additional uncurated samples synthesized by our model using three different backbones: PixArt-$\alpha$, SDv1.5, and SDv2.1, respectively.
Figure 7: Additional qualitative images generated by our PG-SB model with the PixArt-$\alpha$ backbone.
Figure 8: Additional qualitative images generated by our PG-SB model with the SDv1.5 backbone.
Figure 9: Additional qualitative images generated by our PG-SB model with the SDv2.1 backbone.

### v3

@@
-SNOOPI: Supercharged One-step Diffusion Distillation with Proper Guidance
+Supercharged One-step Text-to-Image Diffusion Models with Negative Prompts
 Abstract
-Recent approaches have yielded promising results in distilling multi-step text-to-image diffusion models into one-step ones.
-The state-of-the-art efficient distillation technique, i.e., SwiftBrushv2 (SBv2), even surpasses the teacher model’s performance with limited resources.
-However, our study reveals its instability when handling different diffusion model backbones due to using a fixed guidance scale within the Variational Score Distillation (VSD) loss.
-Another weakness of the existing one-step diffusion models is the missing support for negative prompt guidance, which is crucial in practical image generation.
-This paper presents SNOOPI, a novel framework designed to address these limitations by enhancing the guidance in one-step diffusion models during both training and inference.
-First, we effectively enhance training stability through Proper Guidance - SwiftBrush (PG-SB), which employs a random-scale classifier-free guidance approach.
-By varying the guidance scale of both teacher models, we broaden their output distributions, resulting in a more robust VSD loss that enables SB to perform effectively across diverse backbones while maintaining competitive performance.
-Second, we propose a training-free method called Negative-Away Steer Attention (NASA), which integrates negative prompts into one-step diffusion models via cross-attention to suppress undesired elements in generated images.
-Our experimental results show that our proposed methods significantly improve baseline models across various metrics.
-Remarkably, we achieve an HPSv2 score of 31.08, setting a new state-of-the-art benchmark for one-step diffusion models.
+The escalating demand for real-time image synthesis has driven significant advancements in one-step diffusion models, which inherently offer expedited generation speeds compared to traditional multi-step methods.
+However, this enhanced efficiency is frequently accompanied by a compromise in the controllability of image attributes.
+While negative prompting, typically implemented via classifier-free guidance (CFG), has proven effective for fine-grained control in multi-step models, its application to one-step generators remains largely unaddressed.
+Due to the lack of iterative refinement, as in multi-step diffusion, directly applying CFG to one-step generation leads to blending artifacts and diminished output quality.
+To fill this gap, we introduce Negative-Away Steer Attention (NASA), an efficient method that integrates negative prompts into one-step diffusion models.
+NASA operates within the intermediate representation space by leveraging cross-attention mechanisms to suppress undesired visual attributes.
+This strategy avoids the blending artifacts inherent in output-space guidance and achieves high efficiency, incurring only a minimal 1.89% increase in FLOPs compared to the computational doubling of CFG.
+Furthermore, NASA can be seamlessly integrated into existing timestep distillation frameworks, enhancing the student’s output quality.
+Experimental results demonstrate that NASA substantially improves controllability and output quality, achieving an HPSv2 score of 31.21, setting a new state-of-the-art benchmark for one-step diffusion models.
+Figure 1: Negative prompting are a key technique for improving controllability and image quality in text-to-image diffusion models.
+However, classifier-free guidance (CFG) [19] only supports negative prompt integration in multi-step models due to its iterative nature.
+Our method, NASA, is the first to enable negative prompting in one-step models, allowing for the suppression of unwanted features and providing greater control over image attributes.
+00footnotetext: Equal Contribution
+$\dagger$ Qualcomm AI Research is an initiative of Qualcomm Technologies, Inc.
 1 Introduction
-Diffusion models have recently become popular due to their capacity to produce high-quality, diverse outputs in image [34, 39, 2, 52, 4], 3D [35, 48, 23], audio [17, 10] and video [53, 50, 3, 11] synthesis.
-Unlike other generative models, such as Generative Adversarial Networks (GANs), diffusion models gradually refine their output through a series of steps, allowing them to achieve diverse yet impressively detailed and realistic outputs.
+Diffusion models have recently become popular due to their capacity to produce high-quality, diverse outputs in image [37, 41, 3, 7, 36], 3D [38, 46, 26], audio [21, 15] and video [50, 5] synthesis.
+Unlike other generative models, such as Generative Adversarial Networks (GANs) [16, 23], diffusion models gradually refine their output through a series of steps, allowing them to achieve diverse yet impressively detailed and realistic outputs.
 However, this iterative refinement is both time-consuming and computationally demanding, which limits their practicality in real-world applications.
-As a result, there’s increasing interest in techniques to speed up diffusion models while maintaining output quality.
-Several strategies are being explored to enhance their efficiency, including optimizing model architectures, improving sampling methods, and using latent space representations instead of image space.
-Recently, a particularly promising approach has gained traction: timestep distillation.
-Researchers have found that by adding specific constraints while training another model to mimic a pre-trained multi-step teacher, they can reduce the required steps, creating a faster and more efficient student that retains much of the original quality with minimal adjustments.
-Figure 1: Comparison of various methods, including SNOOPI and alternatives (SwiftBrush, DMD, YOSO, PG-SB), using the PixArt-$\alpha$ backbone.
-Each method generates an image with the prompt ”Vincent van Gogh” (green).
-PG-SB helps stabilize the training process of image-free distillation, addressing the typical instability in such training.
-SNOOPI, a combination of PG-SB and the NASA module, enables effective negative prompting (red) even for one and few-step models, allowing for the removal of specific unwanted features and providing enhanced control over image attributes.
-Current approaches for accelerating text-to-image diffusion models fall into two main categories: image-dependent training and image-free training methods.
-Image-dependent methods, which rely on real images or synthetic ones generated by the teacher models, include adversarial-based approaches [41, 19, 42, 51] and reconstruction-based methods [28, 56, 12, 19].
-While effective, these methods demand huge storage, high memory, extensive computational power and might suffer from mode collapse due to too strong image supervision.
-In contrast, image-free methods [33], which mainly use score distillation losses like Score Distillation Sampling (SDS) [36] or Variational Score Distillation (VSD) [48], avoid the need for image datasets, instead requiring minimal data, often just text prompts.
-Models such as SwiftBrushv2 (SBv2) [7], even when omitting image-regularization term, making it a fully image-free distillation approach, have achieved impressive quality, even outperforming their multi-step counterparts on certain benchmarks.
-Hence, in this paper, we focus on the image-free approach and start with SBv2 as the baseline.
-However, our research identifies two key challenges in current one-step distillation approaches using VSD.
-First, we observe training instability across different architectures, primarily due to the reliance on a fixed guidance scale.
-Second, we identify a broader issue affecting all one-step generators: the lack of negative prompt guidance, which limits the ability to exclude specific elements from generated images—a feature commonly supported in multi-step models [15, 8, 20].
-These issues limit the flexibility and control of one-step models, which reduces their practical effectiveness despite their faster generation times.
-Prior research [54] has attempted to address the stability issues of the VSD framework, but these solutions come with their own limitations.
-Techniques like increased teacher update frequency [54] and adversarial loss terms [42, 19] have been explored to stabilize the training process.
-However, these approaches can add a significant computational burden, partly due to the additional proposed technique itself and partly due to the introduction of image supervision.
-Since we’re focusing on building an image-free distillation, given these limitations, there is a pressing need to conduct a deeper analysis of the VSD framework to understand its challenges and develop more efficient, stable solutions.
-Alongside these stability limitations, as previously mentioned, one-step diffusion models currently lack support for negative prompting, a feature widely used in multi-step diffusion models to enhance image quality and filter out unwanted features.
-In multi-step models, negative prompting techniques are particularly useful for refining alignment with text prompts in text-to-2D generation, and they also help address issues like the Janus problem in text-to-3D generation [8, 20, 1] —where models produce mirrored or duplicated features like faces.
-One common way to integrate negative prompts into multi-step diffusion model sampling is using Classifier-free Guidance (CFG) [15].
-Given two noise predictions from the positive prompt and negative prompt, CFG subtracts negative noise prediction from the positive noise prediction iteratively throughout the denoising process, inherently steering the final output away from the negative prompt.
-However, one-step models do not have the iterative refinement process required by CFG, making it difficult to apply negative prompting in these frameworks.
-This limitation reduces the flexibility and control of one-step models.
-To address this, a novel approach is needed to incorporate the negative prompt signal directly within the one-step generation process, enabling these models to effectively avoid unwanted features without iterative sampling.
-We address these challenges by introducing a robust framework called SNOOPI for the one-step model, illustrated in Fig.
-1, that enhances stability and control during both training and inference.
-(1) We propose Proper Guidance - SwiftBrush (PG-SB), a dynamic guidance approach that varies the guidance scale of teacher models during training to stabilize VSD.
-This flexible CFG improves the adaptability of image-free distillation across diverse diffusion backbones without adding computational or memory overhead.
-(2) We introduce a method called Negative-Away Steer Attention (NASA) used in inference, which incorporates negative prompt guidance into one-step models by adjusting cross-attention layers, effectively reducing unwanted features in generated images.
-By steering attention within intermediate feature space, NASA offers image control capabilities that were previously achievable only in multi-step diffusion models.
-Our contributions are threefold:
-  • We propose Proper Guidance-SwiftBrush (PG-SB), applying a varied guidance scale to teacher model outputs during training, to enhance stability across different diffusion backbones without additional computational costs.
-  • We introduce Negative-Away Steer Attention (NASA), the first method to integrate negative prompt guidance into one-step diffusion models, effectively suppressing unwanted visual attributes.
-  • Extensive experimental results on established benchmarks demonstrate that our method outperforms strong baselines, validating its stability and effectiveness.
+As a result, there is an increasing interest in techniques to speed up diffusion models while maintaining output quality.
+Recent distillation methods [34, 13, 52, 51, 12] compress the entire multi-step generation process into a single step, a breakthrough that enables real-time image synthesis.
+This speedup makes diffusion models practical for applications like content creation and interactive media, where their slow performance was previously a major drawback.
+Despite these impressive speed advancements in one-step generators, we have identified a significant limitation that affects all these models: the inability to support negative prompting [48, 4].
+This feature, which allows users to explicitly exclude unwanted elements, is widely available in multi-step diffusion models but remains notably absent in their single-step alternatives.
+The most widely adopted approach for negative prompting, classifier-free guidance (CFG) [19] works by generating two parallel noise predictions during each denoising step – one incorporating the positive prompt and another using the negative prompt.
+The system then subtracts the negative prediction from the positive one, effectively steering generation away from unwanted attributes.
+This process relies on the iterative nature of multi-step diffusion, where corrections can be applied and refined over dozens of steps.
+One-step generators, however, collapse this entire process into a single forward pass, eliminating the iterative refinement cycles that make CFG possible.
+Without these multiple correction opportunities, there is no straightforward way to incorporate negative guidance into the generation pipeline.
+Consequently, users face an unappealing choice between the instant generation that might include unwanted elements or waiting longer for more precisely controlled results.
+This speed-versus-control dilemma represents the most significant practical barrier to widespread adoption of one-step generators in professional creative workflows, highlighting the urgent need for novel approaches that can bridge this gap.
+To address this issue, we introduce a novel method called Negative-Away Steer Attention (NASA) used in inference, which incorporates negative prompt guidance directly into one-step diffusion models.
+NASA works by adjusting cross-attention layers, effectively reducing unwanted features in generated images.
+By steering attention within intermediate representation space, NASA offers image control capabilities that were previously achievable only in multi-step diffusion models.
+With our newly developed capability, we have explored integrating negative prompting directly into one-step distillation frameworks, focusing on the SwiftBrush (SB) family of models [34, 13].
+The SB approach is compelling for its unique, image-free distillation methodology.
+By matching teacher model probability distributions and applying sophisticated training techniques, SBv2 can outperform its own teacher on standard image synthesis benchmarks.
+Despite these technical achievements, SB models still fall short in human preference evaluations.
+The core challenge lies in the distillation process: without real image supervision, the system lacks clear signals about which generation patterns should be avoided.
+This is precisely where negative prompting and our NASA framework come in.
+Committed to maintaining SB’s valuable image-free distillation approach while incorporating guidance sampling to align with human preference, we provide the critical missing component, a way to steer the model away from undesirable outputs during training.
+NASA enables something previously impossible: guidance sampling within a one-step generator.
+Through this integration, our model achieves an HPSv2 [49] score of 31.21, establishing a new state-of-the-art on human preference metrics for one-step models.
+In summary, our contributions are twofold:
+  • We introduce Negative-Away Steer Attention (NASA), the first method to integrate negative prompt guidance into one-step and few-step diffusion models in both inference and distillation settings.
+  • By integrating NASA into the SwiftBrush (SB) distillation framework, we achieve a new state-of-the-art for one-step diffusion models on human preference metrics.
+This marks the first successful integration of negative prompt guidance directly into one-step model training.
 2 Related Work
-2.1 Text-to-image Diffusion Distillation
-Recent advances in text-to-image diffusion models have focused on accelerating generation through several key methodological streams: conventional distillation, trajectory optimization, and distribution matching.
-Conventional Distillation.
-Early work in model distillation explored two main directions.
-Progressive distillation methods [40, 32] iteratively train a sequence of student models, each halving the required sampling steps while preserving generation quality.
-While effective, this approach necessitates multiple training phases, increasing overall computational cost.
-In parallel, direct distillation techniques [28, 12] generate synthetic noise-image pairs using a teacher model to train students via reconstruction loss, offering a more straightforward training process but often struggling with quality retention.
-Recent work has augmented these methods with adversarial training [19, 42], achieving better image quality at the cost of increased training complexity.
-Trajectory Optimization.
-Trajectory-straightening methods [25, 26] simplify the complex denoising trajectories into near-linear paths, enabling faster sampling without compromising quality.
-Meanwhile, consistency-based frameworks [46, 21, 44] use self-consistency properties to train models to align outputs across multiple points in the ODE-based denoising trajectory.
-Distribution Matching.
-More recent work [33, 7, 55, 54] has explored distribution matching techniques.
+2.1 Distribution Matching Distillation
+Recent work [34, 13, 52, 51] has explored text-to-image diffusion distillation through distribution matching techniques.
 These methods directly align the output distributions of teacher and student models, particularly through Variational Score Distillation (VSD).
 These techniques can be broadly categorized into two approaches: image-dependent training and image-free training.
-Image-dependent methods, which rely on real images for adversarial loss [54, 41] or synthetic ones generated by the teacher models for reconstruction loss [55].
+Image-dependent methods rely on real images for adversarial loss [51, 42] or synthetic ones generated by the teacher models for reconstruction loss [52].
 Despite their effectiveness in preserving image quality, these approaches face significant practical limitations: they require substantial storage capacity, consume high memory during training, demand extensive computational resources, and risk mode collapse due to intensive image supervision.
-In contrast, image-free approaches [33, 7] have demonstrated promising results by relying solely on prompt text for training, offering a significant reduction in resource requirements while maintaining competitive performance.
+In contrast, image-free approaches [34, 13] have demonstrated promising results by relying solely on prompt text for training, significantly reducing resource requirements while maintaining competitive performance.
 However, image-free techniques often encounter training instability, particularly within VSD-based frameworks.
 Existing solutions to stabilize training [52] introduce significant memory and computational overhead, limiting their practical application.
-Our work addresses this gap by introducing a simple yet effective method that enhances training stability without additional computational burden, making image-free variational score distillation more accessible for real-world applications.
-2.2 Sampling Guidance and Negative Prompting
-Classifier-free guidance (CFG) [15] and negative prompting are key techniques for content control in image generation, from text-to-2D to text-to-3D synthesis.
-In multi-step diffusion models, these methods enhance desired image features and suppress undesired ones, improving visual fidelity and reducing artifacts like the Janus problem in 3D tasks [8, 20, 1].
+Our work addresses this gap by introducing a simple yet effective method that enhances training stability without additional computational burden, making image-free variational score distillation produce better models.
+2.2 Sampling Guidance with Negative Prompts
+Classifier-free guidance (CFG) [19] and negative prompting [48, 4] are key techniques for content control during denoising.
+In multi-step diffusion models, these methods enhance desired image features and suppress undesired ones, improving visual fidelity and reducing artifacts [14, 24, 1].
 Score-based approaches like NFSD [24] employ negative prompting to steer generated samples toward real-image distributions, while SDS-Bridge [33] uses negative prompts to model source distributions more accurately.
 Despite their success, negative prompting has yet to be integrated into one-step or few-step diffusion models.
-To bridge this gap, we introduce Negative-Away Steer Attention (NASA), the first method to enable negative prompting in one-step and few-step diffusion models, enhancing control in faster generation settings.
-3 Background
+To bridge this gap, we introduce NASA, the first method to enable negative prompting in one-step and few-step diffusion models, enhancing control in faster generation settings.
+3 Preliminaries
 Diffusion Models [20, 44, 40] produce high-quality images by progressively denoising inputs.
-This technique involves a forward process that gradually adds noise to data and a reverse process that reconstructs the original data by removing noise.
-While the original diffusion models act on image space [16, 45], LDM [38] processes on latent space produced by a pretrained VAE for more efficient computation.
+This technique involves a forward process that gradually adds noise to the data and a reverse process that reconstructs the original data by removing noise.
+While the original diffusion models act on the image space [20, 44], LDM [40] processes on the latent space produced by a pretrained VAE for more efficient computation.
 In particular, starting with a data point $\mathbf{x}_{0}$ sampled from an unknown distribution $q(\mathbf{x}_{0})$, the forward process gradually diffuses $\mathbf{x}_{0}$ into a standard Gaussian noise $\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ through $T$ consecutive timesteps, where $\mathbf{I}$ is the identity matrix.
@@
 Classifier-free Guidance (CFG) [19] is an inference technique designed to enhance the quality of generated images by blending the predictions from both a conditional and an unconditional model.
-At each sampling step, CFG adjusts the denoiser’s output using a control parameter $\kappa>1$, allowing for controlled guidance that aligns more closely with the desired conditions:
+At each sampling step, CFG adjusts the denoiser’s output using a control parameter $\kappa>1$, allowing controlled guidance that aligns more closely with the desired conditions:
 $\hat{\boldsymbol{\epsilon}}_{\phi}(\mathbf{x}_{t},t,{\bf y},\kappa)=(1-\kappa)\,\boldsymbol{\epsilon}_{\phi}(\mathbf{x}_{t},t)+\kappa\,\boldsymbol{\epsilon}_{\phi}(\mathbf{x}_{t},t,{\bf y}).$ (4)
-Negative Prompts provide enhanced control by suppressing unwanted features in the generated content.
+Negative Prompting [48, 4] provides enhanced control by suppressing unwanted features in the generated content.
 Instead of producing an unconditional output, the model generates an output conditioned on the negative prompt $\mathbf{y}^{-}$, as follows:
 $\hat{\boldsymbol{\epsilon}}_{\phi}(\mathbf{x}_{t},t,\mathbf{y},\mathbf{y}^{-},\kappa)=(1-\kappa)\,\boldsymbol{\epsilon}_{\phi}(\mathbf{x}_{t},t,\mathbf{y}^{-})+\kappa\,\boldsymbol{\epsilon}_{\phi}(\mathbf{x}_{t},t,\mathbf{y}).$ (5)
-Variational Score Distillation (VSD) is a powerful framework that utilizes pretrained diffusion-based text-to-image models to enhance text-based generation across both 3D and 2D domains.
-Originally proposed for text-to-3D tasks [48], VSD leverages diffusion-based score matching to align Neural Radiance Fields (NeRFs) with text prompt, enabling the generation of intricate 3D objects [48].
-This approach extends effectively to 2D image synthesis as well, where VSD enables rapid, high-quality text-to-image generation in a single step with models such as SwiftBrush and DMD [33, 7, 55, 54], sidestepping the computational burden of multi-step diffusion methods.
-The central aim of VSD is to ensures that the renderings of a differentiable generator align with the probability density of plausible images as guided by the 2D diffusion model.
-To accomplish this, VSD employs a two-teacher approach that uses a fixed pretrained diffusion model $\boldsymbol{\epsilon}_{\psi}$ and an adaptive LoRA-based teacher $\boldsymbol{\epsilon}_{\phi}$.
-While training, the student model $\boldsymbol{f}_{\theta}$ produces 2D images $\hat{\mathbf{x}}_{0}=\boldsymbol{f}_{\theta}(\mathbf{z},\mathbf{y})$ using an input noise $\mathbf{z}\sim\mathcal{N}(0,I)$ and a text prompt $\mathbf{y}$.
-Noisy images $\hat{\mathbf{x}}_{t}=\alpha_{t}\hat{\mathbf{x}}_{0}+\sigma_{t}\boldsymbol{\epsilon}$ is then fed into both teacher models.
-The LoRA teacher model $\boldsymbol{\epsilon}_{\phi}$ aligns with the student distribution by minimizing a denoising L2 loss on single-step samples.
+Variational Score Distillation (VSD) is a powerful framework that utilizes pretrained diffusion-based text-to-image models to enhance text-based generation across both 3D [46] and 2D [34, 13, 52, 51] domains.
+The central aim of VSD is to align the renderings of a differentiable generator with the probability density of plausible images as guided by the 2D diffusion model.
+To accomplish this, VSD employs a two-teacher approach that uses a fixed pretrained diffusion model $\boldsymbol{\epsilon}_{\psi}$ and an adaptive LoRA-based teacher $\boldsymbol{\epsilon}_{\pi}$.
+While training, the student model $\boldsymbol{f}_{\theta}$ produces 2D images $\hat{\mathbf{x}}_{0}=\boldsymbol{f}_{\theta}(\mathbf{z},\mathbf{y})$ using an input noise $\mathbf{z}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and a text prompt $\mathbf{y}$.
+Noisy images $\hat{\mathbf{x}}_{t}=\alpha_{t}\hat{\mathbf{x}}_{0}+\sigma_{t}\boldsymbol{\epsilon}$ are then fed into both teacher models.
+The LoRA teacher model $\boldsymbol{\epsilon}_{\pi}$ aligns with the student distribution by minimizing a denoising L2 loss on single-step samples.
 This arrangement supports robust and adaptive guidance suitable for a variety of generative architectures.
-The gradient of the learning objective with respect to the student model parameters $\theta$ is defined as:
-$\begin{split}\nabla_{\theta}\mathcal{L}_{\text{VSD}}=\mathbb{E}_{t,\mathbf{y},\mathbf{z},\kappa=C}&\bigg{[}\,w(t)\left(\hat{\boldsymbol{\epsilon}}_{\psi}(\hat{\mathbf{x}}_{t},t,\mathbf{y},\kappa)\right.\\ &\left.-\hat{\boldsymbol{\epsilon}}_{\phi}(\hat{\mathbf{x}}_{t},t,\mathbf{y},\kappa\right)\frac{\partial\boldsymbol{f}_{\theta}(\mathbf{z},\mathbf{y})}{\partial\theta}\bigg{]},\end{split}$ (6)
+The gradient of the learning objective with respect to the student’s parameters $\theta$ is defined as:
+$\begin{split}\nabla_{\theta}\mathcal{L}_{\text{VSD}}=\mathbb{E}_{t,\mathbf{y},\mathbf{z},\kappa=C}&\bigg[\,w(t)\left(\hat{\boldsymbol{\epsilon}}_{\psi}(\hat{\mathbf{x}}_{t},t,\mathbf{y},\kappa)\right.\\ &\left.-\hat{\boldsymbol{\epsilon}}_{\pi}(\hat{\mathbf{x}}_{t},t,\mathbf{y},\kappa\right)\frac{\partial\boldsymbol{f}_{\theta}(\mathbf{z},\mathbf{y})}{\partial\theta}\bigg],\end{split}$ (6)
 where $C$ is a constant and $w(t)$ is the time-dependent weight that adjusts the contribution of each timestep, aligning the student’s outputs with the diffusion model’s predictions.
 Through alternating updates of the student and LoRA teacher, VSD enables efficient, high-fidelity generation across both text-to-image and text-to-3D tasks.
-4 Proper Guidance - SwiftBrush
-Figure 2: FID 2K progression throughout training.
-The default baseline is SwiftBrush for both SDv1.5 and PixArt-$\alpha$.
-DMD2Trick indicates the Two Time-scale Update Rule technique [54], which increases the update frequency of the LoRA teacher to 5 times after each student update.
-These baselines demonstrate notable instability, leading to variability in image quality metrics such as FID.
-In contrast, our approach, which employs CFG with randomly selected values from a uniform distribution, results in a more stable training process and generates higher-quality samples.
+NP Anime Photo CA Paintings Average
+29.62 29.17 28.79 28.69 29.07
+31.66 29.17 30.58 30.38 30.42
+Table 1: The effect of negative prompts (NP) on the HPSv2 score of PixArt-$\alpha$ [9].
+Using generic negative prompts to remove visual artifacts (e.g., “worst quality”, “blurry”, “bad anatomy”) leads to a significant improvement.
+Higher scores indicate better alignment with human preferences.
+4 Observations
 4.1 Motivation
-While the VSD framework offers a solid foundation for model distillation, its training process often experiences instability.
-For example, while successfully working on Stable Diffusion 2.1 (SDv2.1) [38], our re-implementation of SwiftBrush and SwiftBrushv2 (SB) [33, 7] fail in distilling other model backbones such as Stable Diffusion 1.5 (SDv1.5) [38] and PixArt-$\alpha$ [6] when using a fixed CFG scale in the VSD framework.
-Their unstable results are substantiated by quantitative evaluation shown in Fig.
-2.
-Figure 3: Qualitative comparison of different methods for integrating negative prompts into the generation process using a one-step generator.
-As shown, simply applying CFG to a one-step generator is equivalent to blending (i.e., adding) two images together, resulting in an unusable output.
-Alternatively, negating the negative embedding from the positive embedding shows minimal impact on the final image.
-In contrast, NASA is the first method to successfully steer the generation away from negative attributes in a one-step generator, while also producing high-quality results.
-To address this instability, previous studies have proposed techniques such as using image datasets as regularizers, increasing the LoRA teacher update frequency, and adding adversarial loss terms [54, 55].
-However, these methods have two main drawbacks: (1) they rely on image supervision, and (2) they significantly increase memory usage and training time.
-In pursuit of image-free distillation, we instead investigate the existing SB framework to identify the root cause of the instability.
-We begin by examining the impact of the CFG scale in multi-step diffusion models, specifically SDv2.1.
-Our analysis of the MS COCO 2014 benchmark [24], with results shown in Tab.
-1, reveals that model performance shifts considerably as the guidance scale changes.
-Typically, users set the guidance scale based on their specific goal, whether to prioritize image fidelity or text alignment.
-However, this approach is less desirable in training, as we want the model to be versatile across different scenarios.
-Given this, we seek to make the teacher model adaptable across various scenarios without compromising overall performance.
-As shown in Tab.
-1, varying the guidance scale randomly allows us to achieve competitive results across all metrics while maintaining flexibility across different guidance levels.
-$\kappa$ FID $\downarrow$ CLIP $\uparrow$ Precision $\uparrow$ Recall $\uparrow$
-2 10.77 0.31 0.54 0.53
-3 9.60 0.32 0.59 0.50
-4 10.35 0.32 0.60 0.47
-5 11.53 0.33 0.61 0.44
-$\mathcal{U}(2,5)$ 9.86 0.32 0.60 0.48
-Table 1: Results on the zero-shot MS COCO-2014 30K benchmark [24] of SDv2.1 [38] when using PNDM Scheduler with different CFG scale $\kappa$.
-Teacher LoRA Teacher FID $\downarrow$
-10.72
-10.81
-9.51
-9.03
-Table 2: Ablation studies on Random CFG in VSD loss with SDv2.1 backbone on 2K prompts from MS COCO-2014 [24].
-4.2 Method
-Based on our findings, we aim to broaden the guidance distribution provided by the teacher in the SB distillation framework by introducing Proper Guidance - SwiftBrush (PG-SB) which applies a randomized guidance scale during teacher inference.
-Specifically, rather than fixing $\kappa=C$, we sample its value from a uniform distribution $\boldsymbol{\mathcal{U}(\kappa_{\text{min}},\kappa_{\text{max}})}$ where $\kappa_{\text{min}}$ and $\kappa_{\text{max}}$ are predefined constants.
-In the SB distillation pipeline, however, CFG is applied in two places: once to the frozen teacher and once to the LoRA teacher.
-As shown in Tab.
-2, applying this technique only to the LoRA teacher improves model performance.
-However, applying it solely to the frozen teacher destabilizes the distillation process and eventually harms the student performance.
-The gradient of our PG-SB’s VSD loss is defined by:
-$\begin{split}\nabla_{\theta}\mathcal{L}_{\text{VSD}}^{\text{PG-SB}}=&\mathbb{E}_{t,\mathbf{y},\mathbf{z},\boldsymbol{\kappa\sim\mathcal{U}(\kappa_{\text{min}},\kappa_{\text{max}})}}\bigg{[}\,w(t)\left(\hat{\boldsymbol{\epsilon}}_{\psi}(\hat{\mathbf{x}}_{t},t,\mathbf{y},\kappa)\right.\\ &\left.-\hat{\boldsymbol{\epsilon}}_{\phi}(\hat{\mathbf{x}}_{t},t,\mathbf{y},\kappa\right)\frac{\partial\boldsymbol{f}_{\theta}(\mathbf{z},\mathbf{y})}{\partial\theta}\bigg{]}.\end{split}$ (7)
+Figure 2: The inherent single-pass nature of one-step diffusion models renders them incompatible with direct CFG, resulting in artifact-laden outputs.
+Positive prompt is not enough.
+While vision generative models like diffusion models excel at incorporating positive attributes, they fundamentally struggle with understanding negation.
+Consider requesting “A portrait of a friendly monster, not scary” (Fig.
+1), the model typically fails to exclude the specified negative attribute “scary”.
+Negative prompting addresses this limitation by creating a separate, dedicated input for exclusions – completely removing negation terms from the positive prompt and processing them independently.
+This clear separation allows the model to accurately distinguish what to avoid generating, providing significantly better control over the final output.
+Community Need.
+The growing use of negative prompting in the generative AI community further highlights its importance.
+Platforms like Civit AI [10] and various open-source repositories [11, 2, 29] have shown that using both types of prompts together results in more refined, higher-quality model outputs that align more closely with user specifications.
+Users can clearly specify both desired and unwanted attributes in a prompt pair, which is particularly useful when working with models that generate complex or highly detailed content.
+Experimentally, Table 1 and Figure 1 confirm that using negative prompts effectively enhances image quality of the model.
+Critical Gap.
+Despite the established need and widespread adoption of negative prompts in multi-step diffusion models and the rapidly growing prominence of one-step approaches, a significant limitation persists: no existing one-step methods effectively support negative prompting.
+Our proposed method fills this gap by providing a simple, efficient, and training-free mechanism that integrates negative prompt into one-step diffusion models to (1) improve controllability and (2) enhance the quality of the output image.
 Figure 3: Left: An overview of the Negative-Away Steer Attention (NASA) pipeline.
@@
 The final output ($\mathbf{Z}^{\text{NASA}}$) is calculated by subtracting the weighted negative features ($\mathbf{Z}^{-}$) from the positive features ($\mathbf{Z}^{+}$).
-We hypothesize that broadening the output distribution of the LoRA teacher enables it to better align with the student’s output, mainly since the student may not initially perform well at one function evaluation or one NFE generation.
-Interestingly, once the LoRA teacher’s output distribution is expanded, the student can also adapt to the broader range of outputs from the frozen teacher when using a randomized guidance scale, achieving the best results when varied guidance scales are applied to both teachers.
-5 Negative-Away Steer Attention
-5.1 Motivation
-Negative prompting is an essential technique for improving control in text-to-image and text-to-3D synthesis.
-This method helps exclude unwanted elements in generated images and mitigate the Janus problem in 3D generation [20, 31, 1].
-However, no existing approach supports negative prompts in single-step or few-step diffusion models.
-A widely method for implementing negative prompting relies on the CFG mechanism.
+4.2 Challenges
+A widely used method for implementing negative prompting is the CFG mechanism.
 However, applying CFG directly to one-step diffusion models presents significant challenges.
 CFG was designed for multi-step models, where the output is iteratively refined, allowing for gradual adjustments using both negative and unconditional outputs.
-In contrast, one-step models generate images in a single pass without the iterative denoising process, making them inherently less flexible.
-Directly applying CFG to these models leads to an undesirable blending of images, resulting in unnatural artifacts, as shown in Fig.
+In contrast, one-step models generate images in a single pass without the iterative denoising process, making them inherently unsuitable for this method.
+Direct application of CFG to these models leads to undesirable image blending, resulting in unnatural artifacts, as shown in Fig.
+2.
+This highlights a crucial limitation: the CFG mechanism, though effective in multi-step models, is not directly translatable to one-step models without significant modifications.
+5 Proposed Method
+Our proposed methodology is detailed in Sec.
+5.1 for the core mechanism.
+We present two primary implementation strategies for NASA: a training-free approach (NASA–I) and a distillation-based approach (NASA–T), which are presented in Sec.
+5.2 and Sec.
+5.3, respectively.
+5.1 NASA: Negative-Away Steer Attention
+In contrast to CFG implementations that operate within the output space (synthesized image), as commonly practiced in multi-step diffusion models, we propose a novel approach that strategically relocates the guidance mechanism to the intermediate representation space of the diffusion model.
+Specifically, cross-attention layers in these models capture semantic connections between image and text features [17], making them ideal for controlling the feature alignment.
+We hypothesize that cross-attention layer features are particularly well-suited for this approach.
+We selectively manipulate these cross-attention maps to attenuate attention responses associated with negative prompts relative to those elicited by positive prompts, thereby filtering out unwanted semantic attributes.
+This representation-space guidance strategy offers a crucial advantage: it enables the exclusion of undesirable elements prior to the final image synthesis stage.
+This removal mechanism significantly mitigates the risk of generating blending artifacts, which are common byproducts of image-space manipulation techniques.
+Our representation-space guidance also extends to architectures like FLUX [25], which replace traditional cross-attention with joint self-attention blocks.
+[47] show that these blocks still embed distinct image-text cross-attention operations.
+This allows NASA to be seamlessly applied, proving its robustness across different backbones.
+The overall pipeline of the NASA mechanism is illustrated in Fig.
 3.
-This highlights a crucial limitation: the CFG mechanism, though effective in multi-step models, is not directly translatable to one-step models without significant modifications.
-An alternative approach manipulates text embeddings by subtracting the negative embedding from the positive one to steer the model away from undesired attributes.
-However, as shown in Fig.
-3, this method has minimal effect on the generated images, highlighting the need for a more robust solution to negative prompting in one-step models.
-5.2 Method
-Instead of applying CFG in the input space (text embedding) or output space (synthesized image) as in multi-step diffusion, we propose moving CFG to the intermediate feature space of the diffusion model.
-We hypothesize that cross-attention layer features are particularly well-suited for this approach.
-Specifically, cross-attention layers in these models capture semantic connections between image regions and text features [13], making them ideal for controlling feature emphasis.
-By modulating these cross-attention maps, we can filter unwanted features by subtracting attention responses to negative prompts from those to positive prompts.
-This filtering occurs in the feature space, meaning that undesired elements are excluded before the image is generated, reducing the chance of blending artifacts.
-Given the latent’s features at current layer $l$, $\mathbf{Z}_{l}$, and positive text features $\mathbf{c}_{p}$ extracted from CLIP text encoder, we define the query, key, and value matrices as $\mathbf{Q}_{l}=\mathbf{Z}_{l}\mathbf{W}_{q,l}$, $\mathbf{K}_{l}^{+}=\mathbf{c}_{p}\mathbf{W}_{k,l}$, and $\mathbf{V}_{l}^{+}=\mathbf{c}_{p}\mathbf{W}_{v,l}$, where $\mathbf{W}_{q,l}$, $\mathbf{W}_{k,l}$, and $\mathbf{W}_{v,l}$ are projection matrices at current layer $l$.
-For negative text features $\mathbf{c}_{n}$, we use the same projection matrices, defining the key and value matrices as $\mathbf{K}_{l}^{-}=\mathbf{c}_{n}\mathbf{W}_{k,l}$ and $\mathbf{V}_{l}^{-}=\mathbf{c}_{n}\mathbf{W}_{v,l}$.
+Given the latent feature representation at a given layer $l$, denoted as $\mathbf{Z}_{l}$, and positive text features $\mathbf{c}_{p}$ extracted from CLIP text encoder, we define the query, key, and value matrices as $\mathbf{Q}_{l}=\mathbf{Z}_{l}\mathbf{W}_{q,l}$, $\mathbf{K}_{l}^{+}=\mathbf{c}_{p}\mathbf{W}_{k,l}$, and $\mathbf{V}_{l}^{+}=\mathbf{c}_{p}\mathbf{W}_{v,l}$, where $\mathbf{W}_{q,l}$, $\mathbf{W}_{k,l}$, and $\mathbf{W}_{v,l}$ are projection matrices at current layer $l$.
+Analogously, for negative text features $\mathbf{c}_{n}$, we use the same projection matrices to define the corresponding key and value matrices for negative prompt conditioning: $\mathbf{K}_{l}^{-}=\mathbf{c}_{n}\mathbf{W}_{k,l}$ and $\mathbf{V}_{l}^{-}=\mathbf{c}_{n}\mathbf{W}_{v,l}$.
+Method FLUX.1-schnell SDXL-LCM SDXL-DMD2
+4 steps 1 step 4 steps 1 step 4 steps 1 step
+None 23% 44% 43% - 27% 25%
+CFG 30% 0% 14% - 25% 0%
+NASA–I 100% 99% 97% - 100% 100%
+Table 2: Comparison of success rate of unwanted feature removal in generated images. Note that SDXL-LCM [31] does not support one-step inference.
 The NASA mechanism is formulated as follows:
@@
 $\mathbf{Z}^{\text{NASA}}_{l}=\mathbf{Z}_{l}^{+}-\alpha\cdot\mathbf{Z}_{l}^{-},$ (9)
-where $\alpha$ is the scale factor to control the degree of negative feature removal. The overall pipeline of the NASA mechanism is illustrated in Fig. 4.
-6 Experiments
-6.1 Experimental Setup
-Metrics and benchmark.
-In diffusion model evaluation, Fréchet Inception Distance (FID) [14] on the zero-shot MS-COCO2014 benchmark has traditionally been the standard metric.
-However, recent studies [37, 5, 54] suggest that FID does not always correlate with actual image quality and adopt human-focused alignment metrics like Human Preference Score v2 (HPSv2) [49] for benchmarking.
-Our experiments also follow this conclusion, so we choose HPSv2 as the primary metric for evaluating the models.
-Additionally, we use CLIP scores to assess text alignment and precision/recall [22] to evaluate general image quality and mode coverage.
-For completeness, we still report FID results.
-Except for HPSv2, all metrics are computed following the protocols in [18, 7]: generating 30K images using MS-COCO2014 prompts, resizing them to $256\times 256$, and comparing with center-cropped images from the full dataset.
-Implementation details.
-We base our method on SBv2 [7] with our proposed modifications on CFG scale $\kappa$.
-We use total of 3.3M prompts from JourneyDB [47] and a subset of LAION [43].
-All our training is conducted on four NVIDIA A100 40GB GPUs with the total batch size of 64.
-The learning rates are set to $1e^{-6}$ for the student model and $1e^{-3}$ for the LoRA teacher, using the AdamW optimizer [27].
-We utilize three backbones SDv1.5 [38], SDv2.1 [38] and PixArt-$\alpha$ [6] as the frozen teachers and LoRA teachers are initialized with rank $r=64$ and scaling $\gamma=128$.
-The teacher models’ CFG scales are set as follows: $\kappa_{\text{min}}=0.5,\kappa_{\text{max}}=4$ for SDv1.5 and SDv2.1 and $\kappa_{\text{min}}=0.5,\kappa_{\text{max}}=3$ for PixArt-$\alpha$.
-Further details can be found in the Appendix.
-6.2 Evaluation of PG-SB
-Tab.
-4 and Tab.
-5 present a thorough quantitative comparison between our approach and prior distilled text-to-image diffusion models.
-The results consistently show the superiority of our model in HPSv2 scores, with the highest average score across all tested backbones.
-This reinforces the robustness of our method, as discussed in Sec.
-4.2 and demonstrated in Fig.
-2, while also revealing a key insight.
-By gaining a deep understanding of the underlying framework, we made a simple yet impactful change with minimal overhead, leading to significant performance improvements.
-This suggests that integrating complex loss functions may not always be necessary for achieving state-of-the-art results.
-Remarkably, our method achieves these outcomes without requiring image-level supervision, restoring SBv2 as an image-free distillation approach.
-NASA SDXL-LCM SDXL-DMD2 SDXL-DMD2 PG-SB
-4 steps 4 steps 1 step 1 step
-43% 27% 25% 38%
-97% 100% 100% 92%
-Table 3: Comparison of success rate of unwanted feature removal in generated images between models with and without NASA.
-Figure 5: Effect of different scale values (0.0 to 1.0) in NASA with SDXL-DMD2 model, illustrating the progressive influence on visual details and composition.
-Figure 6: Qualitative results of NASA.
-Each row displays images generated with a positive (green) and negative (red) prompt pair.
-Models using NASA effectively exclude unwanted features, aligning outputs with desired traits.
-Method Anime Photo Concept Art Paintings Average
-Stable Diffusion 1.5-based backbone
-SDv1.5 [38] 26.51 27.19 26.06 26.12 26.47
-InstaFlow-0.9B [26] 26.10 26.62 25.92 25.95 26.15
-DMD2 [54] 26.39 27.00 25.80 25.83 26.26
-PG-SB 27.18 27.58 26.69 26.62 27.02
-PG-SB + NASA 27.19 27.59 26.71 26.63 27.03
-Stable Diffusion 2.1-based backbone
-SDv2.1 [38] 27.48 26.89 26.86 27.46 27.17
-SB [33] 26.91 27.21 26.32 26.37 26.70
-SBv2 [7] 27.25 27.62 26.86 26.77 27.13
-PG-SB 27.56 27.84 26.97 27.03 27.35
-PG-SB + NASA 27.71 27.99 27.14 27.27 27.53
-PixArt-$\alpha$-based backbone
-PixArt-$\alpha$ [6] 29.62 29.17 28.79 28.69 29.07
-YOSO [30] 28.79 28.09 28.57 28.55 28.50
-DMD [55] 29.31 28.67 28.46 28.41 28.71
-PG-SB 32.19 29.09 30.39 29.69 30.34
-PG-SB + NASA 32.56 29.55 31.24 30.96 31.08
-Table 4: HPSv2 comparisons between our method and previous works. denotes reported numbers, denotes our rerun based on the publicly available model checkpoints.
-6.3 Evaluation of NASA
-To assess the effectiveness of NASA, we conduct experiments with our model and other one-step and few-step models, such as LCM [29] and DMD2 [54].
-We design a small set of positive and negative prompts and generate 100 images per model, both with and without negative prompts (further details are provided in the Appendix).
+where $\alpha$ is the scale factor to control the degree of negative feature removal.
+5.2 NASA–I: Training-free Setting
+Applicable to One-/Few-Step Models.
+To assess the effectiveness of NASA–I, we carry out experiments with pretrained one-step and few-step models, such as LCM [31], DMD2 [51] and FLUX.1-schnell [25].
+We design a small set of positive and negative prompts and generate 100 images per model, both with and without negative prompts.
 We then calculate the percentage of images that successfully exclude the feature specified by the negative prompt.
@@
 The average results across all prompt sets are summarized in Tab.
-3.
-Our findings show that NASA is compatible with both one-step and few-step models, enabling precise control over unwanted features in generated images.
-Note that the difference in accuracy between PG-SB and other models is primarily attributed to the disparity in backbone size.
-The SDXL-based architecture is well-known for its superior text alignment, which stems from its larger generator backbone and the use of an ensemble of text encoders.
-Additionally, Fig.
+2.
+Our findings show that while traditional CFG is ineffective for one-step and few-step models, NASA–I provides precise control over unwanted features in generated images.
+Additionally, this capability is also qualitatively demonstrated in Fig.
+1, Fig.
 4 and Fig.
-6 present qualitative results of our method.
-Images generated with a conditional negative prompt retain key semantic features such as pose and background, while filtering out unwanted attributes specified by the negative prompt.
-Furthermore, when using negative prompts like “worst quality, low quality, ugly, duplicate, out of frame, deformed, blurry, bad anatomy, watermark”, our method improves image quality, as shown in Tab.
-4 and Tab.
 5.
-Notably, integrating NASA with PG-SB on the PixArt-$\alpha$-based backbone achieves a record-breaking HPSv2 score of 31.08, surpassing all other one-step diffusion models baselines.
-Method #Params NFEs FID $\downarrow$ CLIP $\uparrow$ Precision $\uparrow$ Recall $\uparrow$ Image-free?
+Efficiency Considerations.
+As described in Sec.
+5.1, we reuse the projection matrices of the positive prompt for the negative prompt, ensuring that no additional model parameters are introduced and keeping the model size unchanged.
+Furthermore, applying NASA-I to all one-step and few-step models introduces only a minor computational overhead, increasing FLOPs by just 1.89 %.
+This represents a significant efficiency improvement over CFG, which doubles the computational cost due to the need for separate forward passes for positive and negative prompts.
+Consequently, NASA–I enables negative prompt guidance in one-step diffusion models without introducing substantial latency, making it a practical solution for real-time applications.
+5.3 NASA–T: Distillation Setting
+Given their effectiveness in improving image quality during inference, their potential utility in model distillation is worth exploring.
+In this section, we introduce a method that integrates the negative prompt into one-step diffusion distillation.
+While our approach is applicable to various distillation methods, we select SwiftBrush (SB) [34, 13], an image-free method, as the baseline for improvement due to its efficiency.
+As described in Eq.
+6, SB updates the student model using the gradient of the VSD loss.
+To incorporate the negative prompt $\mathbf{y}^{-}$, we modify the existing formulation by replacing $\hat{\boldsymbol{\epsilon}}_{\psi}(\hat{\mathbf{x}}_{t},t,\mathbf{y},\kappa)$ (see Eq.
+4) with $\hat{\boldsymbol{\epsilon}}_{\psi}(\hat{\mathbf{x}}_{t},t,\mathbf{y},\mathbf{y}^{-},\kappa)$ (see Eq.
+5).
+In this way, the teacher model provides explicit guidance to the student model, helping it learn to steer away from undesirable features.
+Moreover, we incorporate the NASA mechanism into the student model, enabling it to process negative prompts during training.
+This allows the student model to better recognize and suppress unnatural artifacts, leading to improved learning and enhanced image quality.
+Note that with the LoRA teacher model $\hat{\boldsymbol{\epsilon}}_{\pi}$, we can choose whether to incorporate negative prompts.
+Empirically, we found that not using them yields better results.
+Detailed results and discussions are provided in Sec.
+6.3.
+6 Experiments
+Dataset NegOpt HPSv2
+Method CLIP+ $\uparrow$ CLIP- $\downarrow$ Anime $\uparrow$ Photo $\uparrow$ CA $\uparrow$ Paintings $\uparrow$ Average $\uparrow$
+PixArt-$\alpha$-based backbone
+PixArt-$\alpha$ [9] [Teacher] 0.35 0.05 29.62 29.17 28.79 28.69 29.07
+YOSO [32] 0.36 0.08 28.75 28.06 28.52 28.57 28.48
++ NASA–I 0.36 0.06 28.74 28.05 28.56 28.60 28.49 (+0.01)
+DMD [52] 0.35 0.08 29.31 28.67 28.46 28.41 28.71
++ $\text{CFG}=1.5$ 0.34 0.09 30.02 27.07 28.36 28.07 28.38 (-0.33)
++ $\text{CFG}=2.5$ 0.31 0.13 26.74 23.86 25.13 24.66 25.10 (-3.61)
++ NASA–I 0.35 0.05 29.33 28.71 28.49 28.53 28.77 (+0.06)
+SBv2 0.36 0.09 32.19 29.09 30.39 29.69 30.34
++ NASA–I 0.36 0.06 32.60 29.58 31.09 30.65 30.98 (+0.64)
++ NASA–T 0.35 0.08 32.33 29.26 30.75 30.10 30.61 (+0.27)
++ NASA–T + $\text{CFG}=1.5$ 0.34 0.10 29.47 26.50 28.22 27.68 27.97 (-2.37)
++ NASA–T + NASA–I 0.35 0.06 32.66 29.67 31.31 30.71 31.09 (+0.75)
++ NASA–T + NASA–I ($\alpha=1$) 0.35 0.05 32.65 29.65 31.45 31.06 31.21 (+0.87)
 Stable Diffusion 1.5-based backbone
-SDv1.5 ($\textit{cfg}=3$) [38] 0.9B 25 8.78 0.30 0.59 0.53
-UFOGen [51] 0.9B 1 12.78 - - -
-InstaFlow-0.9B [26] 0.9B 1 13.33 0.30 0.53 0.45
-DMD [55] 0.9B 1 11.49 0.32 - -
-DMD2 [54] 0.9B 1 8.29 0.30 0.59 0.52
-PG-SB 0.9B 1 10.08 0.31 0.57 0.47
-PG-SB + NASA 0.9B 1 9.94 0.31 0.57 0.48
+SDv1.5 [40] [Teacher] 0.31 0.06 26.51 27.19 26.06 26.12 26.47
+InstaFlow-0.9B [28] 0.33 0.11 26.10 26.62 25.92 25.95 26.15
++ NASA–I 0.33 0.10 26.15 26.68 25.93 25.98 26.19 (+0.04)
+DMD2 [51] 0.31 0.10 26.39 27.00 25.80 25.82 26.25
++ NASA–I 0.31 0.09 26.39 27.02 25.80 25.83 26.26 (+0.01)
+SBv2 0.32 0.09 27.18 27.58 26.69 26.62 27.02
++ NASA–I 0.32 0.08 27.19 27.59 26.71 26.63 27.03 (+0.01)
++ NASA–T 0.35 0.10 27.64 27.94 27.19 27.03 27.45 (+0.43)
++ NASA–T + NASA–I 0.35 0.09 27.65 27.97 27.18 27.04 27.46 (+0.44)
 Stable Diffusion 2.1-based backbone
-SDv2.1 ($\textit{cfg}=3$) [38] 0.9B 25 9.60 0.32 0.59 0.50
-SB [33] 0.9B 1 15.46 0.30 0.47 0.46
-SBv2 [7] 0.9B 1 8.14 0.32 0.57 0.52
-PG-SB 0.9B 1 8.91 0.34 0.57 0.49
-PG-SB + NASA 0.9B 1 8.83 0.34 0.57 0.50
-PixArt-$\alpha$-based backbone
-PixArt-$\alpha$ ($\textit{cfg}=4.5$) [6] 0.6B 20 26.85 0.32 0.52 0.23
-YOSO [30] 0.6B 1 26.04 0.30 0.46 0.30
-DMD [55] 0.6B 1 30.22 0.32 0.55 0.18
-PG-SB 0.6B 1 22.03 0.33 0.52 0.24
-PG-SB + NASA 0.6B 1 21.98 0.33 0.52 0.25
-Table 5: Quantitative comparisons between our method and others on zero-shot MS COCO-2014 benchmark.
-For multi-step SD models, we report each with the CFG scale that returns the best FID.
-denotes reported numbers, denotes our rerun based on the publicly available model checkpoints.
-‘-’ denotes unreported results.
+SDv2.1 [40] [Teacher] 0.33 0.06 27.48 26.89 26.86 27.46 27.17
+SBv2 [13] 0.34 0.10 27.25 27.62 26.86 26.77 27.13
++ NASA–I 0.35 0.08 27.44 27.84 26.91 27.02 27.30 (+0.17)
+SBv2 0.37 0.10 27.56 27.84 26.97 27.03 27.35
++ NASA–I 0.37 0.08 27.70 28.00 27.12 27.22 27.51 (+0.16)
++ NASA–T 0.36 0.09 28.00 28.65 27.44 27.26 27.84 (+0.49)
++ NASA–T + NASA–I 0.36 0.07 28.04 28.68 27.51 27.50 27.93 (+0.58)
+Table 3: Quantitative comparisons between our method and previous works.
+NASA–I refers to our training-free method introduced in Sec.
+5.2, while NASA–T denotes our distillation-based approach described in Sec.
+5.3.
+SBv2 refers to our reimplementation of SBv2 with a randomly sampled CFG scale $\kappa$.
+Figure 4: Effect of different scale values (0.0 to 1.0) in NASA-I with SDXL-DMD2 [51] for both one-step and few-step settings, illustrating the progressive influence on visual details and composition.
+Figure 5: Qualitative results of NASA-I in FLUX.1-schnell [25].
+Figure 6: Qualitative results of NASA-I for text-to-video generation in CausVid [53].
+Method VBench-Long benchmark
+Aesthetic Quality $\uparrow$ Imaging Quality $\uparrow$
+None 61.98 67.12
+NASA–I 63.33 67.36
+Table 4: CausVid quality metrics.
+6.1 Experimental Setup
+Metrics and Benchmark.
+In diffusion model evaluation, Fréchet Inception Distance (FID) [18] on the zero-shot MS-COCO2014 [27] benchmark has been the standard metric.
+However, recent studies [39, 8, 51] suggest that FID in this benchmark does not always correlate with actual image quality and adopt human-focused alignment metrics like Human Preference Score v2 (HPSv2) [49] for benchmarking.
+Our experiments also follow this conclusion, so we choose HPSv2 as the primary metric for evaluating the models.
+Moreover, due to the absence of negative prompts in the MS-COCO2014 benchmark, we utilize the NegOpt dataset from [35], which contains pairs of positive and negative prompts.
+From this dataset, we randomly sample 30K prompt pairs to construct a subset for evaluation.
+Text alignment is then assessed using CLIP scores, where CLIP+ evaluates alignment with the positive prompt and CLIP- evaluates alignment with the negative prompt.
+Training Datasets.
+Following SBv2 [13], we use total of 3.3M prompts from JourneyDB [45] and a subset of LAION [43] [^1] All datasets were downloaded and evaluated at MovianAI..
+For negative prompts in the distillation setting (Sec.
+5.3), we construct a dataset of 10K negative prompts derived from commonly used terms like “bad photo, duplicate, low resolution, low quality”.
+Implementation Details.
+We base our method on SBv2 [13] with our proposed modifications in Sec.
+5.3.
+We utilize three backbones SDv1.5 [40], SDv2.1 [40] and PixArt-$\alpha$ [9] as the frozen teachers and LoRA teachers are initialized with rank $r=64$ and scaling $\gamma=128$.
+The learning rates are set to $1e^{-6}$ for the student model and $1e^{-3}$ for the LoRA teacher, using the AdamW optimizer [30].
+All our training is conducted on four NVIDIA A100 40GB GPUs with the total batch size of 64 for the SD models and 32 for PixArt-$\alpha$.
+Moreover, [6] found that instead of using a fixed CFG scale $\kappa$ during training, sampling $\kappa$ from a uniform distribution within a given range improves training robustness.
+Based on this insight, we set the teacher models’ CFG scales as follows: $\kappa\sim\mathcal{U}(0.5,4)$ for SDv1.5 and SDv2.1, and $\kappa\sim\mathcal{U}(0.5,3)$ for PixArt-$\alpha$.
+For NASA–T, in each iteration, we uniformly sample the scale $\alpha\sim\mathcal{U}(0,1)$ to help the student model adapt to different conditions and enhance generalization.
+For NASA–I, unless explicitly specified, we set $\alpha=0.1$ for SDv1.5-based models, $\alpha=0.2$ for SDv2.1-based models, and $\alpha=0.5$ for PixArt-$\alpha$-based models.
+Further details can be found in the Appendix.
+NASA
+Teacher
+LoRA Teacher
+HPSv2 Avg. $\uparrow$ 30.34 29.84 29.80 30.61 29.89
+Table 5: Ablation studies on the integration of negative prompts in model distillation using the PixArt-$\alpha$ backbone.
+Figure 7: Ablation study on the NASA scale using our distilled model based on the PixArt-$\alpha$ backbone.
+6.2 Main Results
+Quantitative Results.
+Table 3 presents a comparison between our approach and prior distilled text-to-image diffusion models.
+Using our reimplementation of SBv2 [13] as the baseline, NASA–T consistently enhances image quality across all tested backbones.
+Compared to other methods, our model achieves the highest HPSv2 scores on average, demonstrating its clear superiority in image generation quality.
+This suggests that integrating complex loss functions like DMD2 [51] may not always be necessary for achieving state-of-the-art results.
+Moreover, while applying CFG directly degrades one-step model outputs, NASA–I further enhances image quality across all models when using negative prompts like “worst quality, low quality, ugly, duplicate, out of frame, deformed, blurry, bad anatomy, watermark”.
+Notably, our method on the PixArt-$\alpha$-based backbone achieves a record-breaking HPSv2 score of 31.21, surpassing all other one-step diffusion models baselines.
+Regarding the NegOpt dataset, with the support of NASA–I, all models demonstrate a reduction in the CLIP- score while maintaining the CLIP+ score.
+This indicates our method effectively eliminates unwanted features in generated outputs while preserving desired traits.
+Qualitative Results.
+As discussed in Sec.
+5.2, NASA–I is a versatile, training-free method for both one-step and few-step diffusion models.
+Figure 4 and Figure 5 present qualitative results of integrating NASA–I into the SDXL-DMD2 [51] and FLUX.1-schnell [25] models, respectively.
+When using a conditional negative prompt, the generated images retain key semantic features while effectively suppressing unwanted attributes specified by the negative prompt.
+Application to Text-to-Video Models.
+To demonstrate the versatility of our approach, we extended NASA-I to a few-step video diffusion model, CausVid [53].
+As shown in Fig.
+6, our method effectively suppresses unwanted visual attributes while preserving key semantic content and temporal consistency.
+Quantitative evaluation on the VBench-Long [22] benchmark further validates this.
+Table 4 shows that using NASA–I with common negative prompts like “low quality, bad anatomy, blurry, distortion, ugly, low resolution, unclear details” improves both the aesthetic and imaging quality scores of the generated videos.
+This highlights NASA–I’s potential as a general-purpose tool for enhancing guided generation across different modalities.
+6.3 Ablation Studies
+Table 5 presents an ablation on integrating negative prompts during model distillation with the PixArt-$\alpha$ [9] backbone.
+The baseline is our reimplementation of SBv2 with a randomly sampled CFG scale $\kappa$.
+Simply applying negative prompts only in the teacher and/or the LoRA teacher models slightly degrades the HPSv2 score, as the student model is not explicitly trained on this negative conditioning.
+In contrast, incorporating NASA with negative prompts during training allows the student model to better capture and utilize this conditioning, leading to improved performance.
+Furthermore, we find that integrating negative prompts solely into the teacher model yields better results than applying them to both the teacher and LoRA teacher models.
+Using our PixArt-$\alpha$-based distilled model as the base, we varied the NASA scale $\alpha$ and evaluated the results with HPSv2 on the benchmark dataset.
+Figure 7 demonstrates that NASA–I generally enhances image synthesis, resulting in outputs that better align with human preferences.
+However, if the NASA scale $\alpha$ is set too high, the negative attributes may overly dominate the positive attributes, adversely affecting the final image quality.
 7 Discussion and Conclusion
-Limitations: Despite its promising results, PG-SB currently lacks support for few-step models, limiting its applicability to downstream tasks.
-Additionally, while NASA proves effective, achieving high-quality images depends on selecting an appropriate scale for removing negative features, though the parameter does not require extensive tuning.
-Furthermore, NASA is designed specifically for architectures with cross-attention layers, leaving support for other architectures, such as MM-DiT [9], as future work.
-Conclusion: This paper introduces SNOOPI, a framework that improves stability and control in one-step diffusion models for both training and inference.
-PG-SB dynamically adjusts the guidance scale to stabilize VSD across various model backbones without increasing computational cost.
-Additionally, NASA is the first approach to incorporate negative prompt guidance in one-step diffusion models, enabling the effective reduction of unwanted features in generated images through cross-attention adjustments.
-Experimental results demonstrate that SNOOPI outperforms strong baselines, showcasing its stability and effectiveness.
+This paper introduces NASA, the first method to integrate negative prompt guidance into one-step and few-step diffusion models for both inference and distillation.
+By leveraging cross-attention adjustments, NASA effectively suppresses unwanted features in generated images.
+Experimental results show that NASA enhances image quality and outperforms strong baselines, establishing a new state-of-the-art in one-step model distillation based on human preference metrics.
+While NASA’s effectiveness relies on selecting an appropriate scale for suppressing negative features, this parameter does not require extensive tuning.
 Supplementary Material
 8 Implementation Details
-PG-SB: Following SBv2 [7], we also use the clamped CLIP loss with a margin of $\tau=0.37$, starting with a weight of 0.1 and gradually reducing to zero.
-Tab.
-6 provides the additional hyperparameters for training PG-SB.
-Hyperparameter SDv1.5 SDv2.1 PixArt-$\alpha$
-Dataset JDB + LAION JDB + LAION JDB
-Batch size 64 64 64
-Training iterations 60k 40k 60k
-Mixed-Precision (BF16) Yes Yes Yes
-$(\kappa_{\text{min}},\kappa_{\text{max})}$ (0.5, 4) (0.5, 4) (0.5, 3)
-Clip weight 0.1 0.1 0.1
-$\tau$ 0.37 0.37 0.37
-$lr$ of student $1e^{-6}$ $1e^{-6}$ $1e^{-6}$
-$lr$ of LoRA teacher $1e^{-3}$ $1e^{-3}$ $1e^{-3}$
-LoRA rank $r$ 64 64 64
-LoRA scaling $\gamma$ 128 128 128
-Table 6: Hyperparameters used for training PG-SB.
-Positive prompt Negative prompt SDXL-LCM SDXL-DMD2 SDXL-DMD2 PG-SB
-4 steps 4 steps 1 step 1 step
-“A photo of a person” “male” 52% $\rightarrow$ 99% 26% $\rightarrow$ 100% 29% $\rightarrow$ 100% 42% $\rightarrow$ 81%
-“female” 48% $\rightarrow$ 100% 74% $\rightarrow$ 100% 71% $\rightarrow$ 100% 58% $\rightarrow$ 88%
-“A photo of a person” “young” 55% $\rightarrow$ 100% 44% $\rightarrow$ 100% 7% $\rightarrow$ 100% 39% $\rightarrow$ 96%
-“old” 45% $\rightarrow$ 92% 56% $\rightarrow$ 100% 93% $\rightarrow$ 100% 61% $\rightarrow$ 79%
-“A photo of a pet” “cat” 36% $\rightarrow$ 100% 12% $\rightarrow$ 100% 60% $\rightarrow$ 100% 32% $\rightarrow$ 99%
-“dog” 64% $\rightarrow$ 100% 88% $\rightarrow$ 100% 40% $\rightarrow$ 100% 68% $\rightarrow$ 99%
-Table 7: Comparison of success rates of unwanted feature removal in generated images before and after applying the NASA model.
-Method Anime Photo Concept Art Paintings Average
-Stable Diffusion 1.5-based backbone
-InstaFlow-0.9B [26] 26.10 26.62 25.92 25.95 26.15
-InstaFlow-0.9B + NASA ($\alpha=0.19$) 26.24 26.74 26.04 26.00 26.26 (+0.11)
-DMD2 [54] 26.39 27.00 25.80 25.83 26.26
-DMD2 + NASA ($\alpha=0.04$) 26.41 27.02 25.80 25.83 26.27 (+0.01)
-PG-SB 27.18 27.58 26.69 26.62 27.02
-PG-SB + NASA ($\alpha=0.04$) 27.19 27.59 26.71 26.63 27.03 (+0.01)
-Stable Diffusion 2.1-based backbone
-SBv2 [7] 27.25 27.62 26.86 26.77 27.13
-SBv2 + NASA ($\alpha=0.30$) 27.45 27.85 26.93 27.09 27.33 (+0.20)
-PG-SB 27.56 27.84 26.97 27.03 27.35
-PG-SB + NASA ($\alpha=0.30$) 27.71 27.99 27.14 27.27 27.53 (+0.18)
-PixArt-$\alpha$-based backbone
-YOSO [30] 28.79 28.09 28.57 28.55 28.50
-YOSO + NASA ($\alpha=0.20$) 28.80 28.10 28.62 28.57 28.52 (+0.02)
-DMD [55] 29.31 28.67 28.46 28.41 28.71
-DMD + NASA ($\alpha=0.45$) 29.34 28.71 28.50 28.52 28.77 (+0.06)
-PG-SB 32.19 29.09 30.39 29.69 30.34
-PG-SB + NASA ($\alpha=0.80$) 32.56 29.55 31.24 30.96 31.08 (+0.74)
-Table 8: HPSv2 comparisons between our method and previous works. denotes reported numbers, denotes our rerun based on the publicly available model checkpoints.
-NASA: We provide the PyTorch pseudo-code for the NASA Attention algorithm, as outlined in Sec. 5. The implementation is straightforward and is provided in Algorithm 1.
-Algorithm 1 NASA Attention.
-class NASA_AttnProcessor(nn.Module):
-def __init__(self, scale=1.0):
+NASA. We provide the PyTorch pseudo-code for the NASA algorithm, as outlined in Section 5.1 in the main paper. The implementation is straightforward and is provided in Algorithm 1.
+Algorithm 1 NASA.
+class NASA_AttnProcessor($n$.Module):
+def __init__(self, nasa_scale=1.0):
 super().__init__()
-self.scale = scale
+self.nasa_scale = nasa_scale
 def __call__(
-self, attn, z, emb, neg_emb, attn_mask, neg_attn_mask, temb, scale=1., *args, **kwargs,
+self, attn, z, emb, neg_emb, attn_mask, neg_attn_mask, temb, *args, **kwargs,
 # Input preparation
@@
 # Equation 10 in main paper
-z_nasa = z - self.scale * neg_z
+z_nasa = z - self.nasa_scale * neg_z
 z_nasa = attn.to_out[0](z_nasa)
@@
 return z_nasa
-9 More Analysis on NASA
-9.1 Experimental Details in Tab. 3
-For clarification, in Tab.
-3, we evaluate the effectiveness of NASA across four models: SDXL-LCM, SDXL-DMD2 (with both 1-step and 4-step settings), and our PG-SB model using the SDv2.1 backbone.
-For the “without NASA” case, all models use a scale of $\alpha=0$, while for the “with NASA” case, a scale of $\alpha=0.5$ is applied.
-The evaluation involves 6 positive-negative prompt pairs.
-For each model and prompt pair, 100 images are generated.
-Human evaluations are then used to determine the percentage of images that successfully exclude the feature specified by the negative prompt.
-The detailed results from Tab.
-3 are provided in Tab.
-7.
-9.2 More Quantitative Results
-Tab.
-8 presents additional quantitative results when applying NASA to various models.
-Using a negative prompt such as “worst quality, low quality, ugly, duplicate, out of frame, deformed, blurry, bad anatomy, watermark,” our NASA method demonstrates a slight improvement in the HPSv2 scores across all models.
-10 Additional Qualitative Results of PG-SB
-Fig. 7, Fig. 8, and Fig. 9 present additional uncurated samples synthesized by our model using three different backbones: PixArt-$\alpha$, SDv1.5, and SDv2.1, respectively.
-Figure 7: Additional qualitative images generated by our PG-SB model with the PixArt-$\alpha$ backbone.
-Figure 8: Additional qualitative images generated by our PG-SB model with the SDv1.5 backbone.
-Figure 9: Additional qualitative images generated by our PG-SB model with the SDv2.1 backbone.
+NASA–T.
+Following SBv2 [13], we also use the clamped CLIP loss with a margin of $\tau=0.37$, starting with a weight of 0.1 and gradually reducing to zero.
+Table 6 provides the additional hyperparameters for training SBv2 with NASAT.
+Hyperparameter SDv1.5 SDv2.1 PixArt-$\alpha$
+Dataset JDB + LAION JDB + LAION JDB
+Batch size 64 64 32
+Training iterations 60k 40k 50k
+Mixed-Precision (BF16) Yes Yes Yes
+$\kappa$ $\mathcal{U}(0.5,4)$ $\mathcal{U}(0.5,4)$ $\mathcal{U}(0.5,3)$
+$\alpha$ $\mathcal{U}(0,1)$ $\mathcal{U}(0,1)$ $\mathcal{U}(0,1)$
+Clip weight 0.1 0.1 0.1
+$\tau$ 0.37 0.37 0.37
+lrlr of student $1e^{-6}$ $1e^{-6}$ $1e^{-6}$
+lrlr of LoRA teacher $1e^{-3}$ $1e^{-3}$ $1e^{-3}$
+LoRA rank $r$ 64 64 64
+LoRA scaling $\gamma$ 128 128 128
+Table 6: Hyperparameters used for training SBv2 [13] with NASA–T.
+9 More Qualitative Results
+Figure 8 and Figure 9 illustrate the visual effect of different scale values $\alpha$ ranging from 0.0 to 1.0 in the NASA method when being applied on SDXL-DMD2.
+A higher scale value $\alpha$ results in more effective removal of the feature specified by the negative prompt.
+Figure 8: Additional qualitative images of applying NASA–I to SDXL-DMD2 1-step.
+Figure 9: Additional qualitative images of applying NASA–I to SDXL-DMD2 4-step.